### v1

# A Debiased Estimator for the Mediation Functional in Ultra-High-Dimensional Setting in the Presence of Interaction Effects

Abstract
Mediation analysis is crucial in many fields of science for understanding the mechanisms or processes through which an independent variable affects an outcome, thereby providing deeper insights into causal relationships and improving intervention strategies. Despite advances in analyzing the mediation effect with fixed/low-dimensional mediators and covariates, our understanding of estimation and inference of mediation functional in the presence of (ultra)-high-dimensional mediators and covariates is still limited. In this paper, we present an estimator for mediation functional in a high-dimensional setting that accommodates the interaction between covariates and treatment in generating mediators, as well as interactions between both covariates and treatment and mediators and treatment in generating the response. We demonstrate that our estimator is $\sqrt{n}$-consistent and asymptotically normal, thus enabling reliable inference on direct and indirect treatment effects with asymptotically valid confidence intervals. A key technical contribution of our work is to develop a multi-step debiasing technique, which may also be valuable in other statistical settings with similar structural complexities where accurate estimation depends on debiasing. Corresponding authors.
Keywords— Causal Mediation Analysis; Debiased Estimation; Ultra-High-Dimensional Models; Interaction Effects
1 Introduction
The most commonly targeted quantity in the field of causal inference is the total causal effect, which captures the causal effect of a treatment, action, or policy on an outcome variable of interest. Causal mediation analysis goes one step further and delves into identifying the mechanisms through which treatment influences the outcome variable. Hence, it plays a crucial role in deepening our understanding of the intricate causal relationships within the variables of the system under study. Mediation analysis has widespread use in many fields of science, including social science, behavioral science, economics, decision-making, epidemiology, and neuroscience. In the past three decades, the topic has also received much attention in the statistical literature (see, e.g., (Albert, , 2008; Lindquist, , 2012; Baron and Kenny, , 1986; Richiardi et al., , 2013; Ten Have and Joffe, , 2012)).
The advancement of technology aids in the collection of an extensive array of variables in various disciplines such as brain imaging, genetics, epidemiology, and public health studies. Consequently, many scientific domains are faced with scenarios in which mediation analysis must be carried out in a high-dimensional setting; that is, the number of potential mediators is quite large and can even be greater than the number of observations. This challenge is particularly pronounced in environmental epigenetics when using omics data. However, most existing methodologies rely on fixed-dimensional mediators, which cannot be applied directly to high-dimensional data. This has motivated recent work on developing methodologies for high-dimensional mediation analysis. Several methodologies have been developed for estimating direct and indirect treatment effects in the presence of high-dimensional mediators. Zhang et al., (2016) and Perera et al., (2022) proposed to first screen mediators that are strongly associated with outcome and run standard analysis on these chosen low-dimensional mediators. Huang and Pan, (2016) and Chén et al., (2018) suggested reducing dimension by a general mapping to low-dimensional spaces instead of variable selection. Zhang et al., (2018) also performed dimension reduction but based on using distance metrics and grouping individually weak signals. Zhao and Luo, (2016) proposed a pathwise Lasso method for selecting sparse mediation pathways in the presence of high-dimensional mediators. None of the aforementioned works presented any theoretical study for their proposed methodology. However, limited theoretical progress has also been made in the literature. For example, Guo et al., (2022) provides an estimator (along with theoretical guarantees) of direct and indirect treatment effects in the presence of high-dimensional mediators and fixed-dimensional observed confounders and exposures. Guo et al., (2023) considered testing the effect the high-dimensional mediators on the response variable in the presence of fixed dimensional exposures. Guo et al., (2024) expanded the estimation method of Guo et al., (2023) by incorporating a general link function to characterize the relationship between the response variable and the mediators and exposures. Lin et al., (2023) proposed a methodology to estimate the indirect effect in the presence of multivariate mediators and exposures. Their method allows both the dimension of the mediators and the exposures to grow with the sample size. A recent study by Rakshit and Guo, (2024) introduced an estimator for the direct (and indirect) treatment effect, accommodating high-dimensional mediators and their interaction with treatment. Our approach generalizes this work within a linear model framework, extending it to include both high-dimensional covariates, mediators, and their interaction with the treatment indicator. We note that all these papers either consider no observed covariates or fixed-dimensional observed covariates in the setting, i.e., they do not allow both the mediators and the observed covariates to be (ultra) high-dimensional. Moreover, they do not allow for interaction between exposures and mediators or between exposures and the observed covariates.
Our contribution: In this paper, we propose a novel methodology for estimating the mediation functional in the presence of ultra-high-dimensional (UHD) covariates and mediators (where the dimension can be exponentially large compared to the sample size). We consider the case where the response variable is generated from a UHD linear model (with respect to both the covariates and the mediators), and the UHD mediators themselves are generated from a linear model with respect to the UHD covariates. Unlike the majority of the existing literature on high-dimensional mediation analysis, our setup allows for interaction with binary exposure for both data-generating processes. To the best of our knowledge, this is the first work to provide a methodology for estimating direct and indirect treatment effects in the presence of both UHD covariates and mediators, along with their interaction with a binary exposure variable. Our method of estimating mediation functional essentially consists of three high-dimensional regressions. We first regress response on both covariates and mediators using treatment observations and then regress mediators on covariates using control observations. Since both regressions are high-dimensional, we carefully use a debiasing technique to remove the bias, which necessitates the third high-dimensional regression. We demonstrate that, under fairly general assumptions, our proposed estimator of mediation functional is $\sqrt{n}$-consistent and asymptotically normal (where $n$ is the sample size). We also provide a consistent estimator for the asymptotic variance, which aids in constructing an asymptotically valid confidence interval.
Our debiasing approach is primarily motivated by the large body of literature on the debiasing techniques for estimating a contrast of high-dimensional parameters in a linear model (see, e.g., (Javanmard and Montanari, , 2014; Van de Geer et al., , 2014; Zhang and Zhang, , 2014)). The use of the debiasing technique in causal inference was also recently employed by Athey et al., (2018), where the authors developed a methodology to estimate the total effect of a binary treatment on an outcome in the presence of high-dimensional covariates. We generalize that work by estimating the direct and indirect parts of the causal effect separately while allowing the mediator to be UHD and have interactions with the treatment variable. Our proposed method can be considered as a high-dimensional counterpart of the influence function-based approach of Tchetgen Tchetgen and Shpitser, (2012). In particular, we demonstrate that our estimator possesses a second-order bias with product form, similar to the influence function-based approach for the fixed-dimensional setting, as established by Ghassami, (2024) based on the influence function of Tchetgen Tchetgen and Shpitser, (2012).
The rest of the paper is organized as follows. After introducing the problem setting in Section 2, we describe our debiased estimation procedure in Section 3. In Section 4, we provide the bias analysis of our estimator and its connection to the fixed-dimensional setting. In Section 5, we provide a theoretical justification for our approach under high-dimensional asymptotic. The proof of the main theorem, along with technical lemmas, can be found in the Appendix. We conclude in Section 6 and discuss future work.
2 Notations and Problem Setting
For a random variable/vector $X$, we denote $\mathbb{E}[X]$ and $\Sigma_{X}$ as the expectation and the variance/covariance matrix of $X$, respectively. For a matrix $\mathbf{X}$, we use $\mathbf{X}_{i,.}$ (respectively $\mathbf{X}_{.,i}$) to denote its $i^{th}$ row (respectively $i^{th}$ column). For a matrix $\mathbf{X}$, $\|\mathbf{X}\|_{1}$, $\|\mathbf{X}\|_{2}$, $\|\mathbf{X}\|_{\infty}=\max_{i,j}|\mathbf{X}_{ij}|$, $\|\mathbf{X}\|_{1,\infty}=\underset{1\leq i\leq m}{\max}\|\mathbf{X}_{i,\cdot}\|_{1}$, $\|\mathbf{X}\|_{\text{op}}=\underset{\|v\|=1}{\sup}\|\mathbf{X}v\|$ and $\|\mathbf{X}\|_{F}=\sqrt{\sum_{i,j}X^{2}_{i,j}}$ represent its $L_{1}$-norm, $L_{2}$-norm, $L_{\infty}$-norm, $L_{1,\infty}$-norm, operator norm and Frobenius norm, respectively. $\mathbf{1}_{q}$ and $\mathbf{0}_{q}$ are the vectors consisting of all ones and all zeros, respectively in $\mathbb{R}^{q}$. $I_{n}$ represents the identity matrix of size $n$. $\mathcal{N}(\mu,\sigma^{2})$ represents the Gaussian distribution with mean $\mu$ and variance $\sigma^{2}$. We denote maximum and minimum by $a\vee b=\max\{a,b\}$ and $a\wedge b=\min\{a,b\}$.
We consider a setting with a binary treatment variable $A\in\{0,1\}$ (also called exposure or action variable), a set of pre-treatment covariates $X=[X_{1}\cdots X_{p}]^{\top}\in\mathbb{R}^{p}$, an outcome variable of interest, denoted by $Y$, and a set of post-treatment, pre-outcome variables $M=[M_{1}\cdots M_{q}]^{\top}\in\mathbb{R}^{q}$. We consider the challenging setting where $p\wedge q>>n$. Using the potential outcome notations (Rubin, , 1974), let $Y^{(a,m)}$ denote the potential outcome of $Y$, had the treatment and mediator variables been set to values $A=a$ and $M=m$. Similarly, we define $M^{(a)}$ as the potential outcome of $M$ had the treatment variables been set to the value $A=a$. Based on variables $Y^{(a,m)}$ and $M^{(a)}$, we define $Y^{(a)}=Y^{(a,M^{(a)})}$ and $Y^{(m)}=Y^{(A,m)}$. An iid sample $(X_{i},A_{i},M_{i},Y_{i})_{i=1}^{n}$ of size $n$ is given, where we assume
  • $Y=Y^{(a,m)}$ if $A=a$, $M=m$, and
  • $M=M^{(a)}$ if $A=a$,
which is referred to as the consistency assumption, requiring that the realized outcome is equal to the potential outcome corresponding to the observed treatment and mediator, and the realized mediator is equal to the potential mediator corresponding to the observed treatment. We also assume that the probability of receiving treatment and control and any value of the mediator are bounded away from zero, i.e., for some $e>0$, we have
  • $0<e<\mathbb{P}(a\mid x)$, for all $a,x$, and
  • $0<e<\mathbb{P}(m\mid a,x)$, for all $a,x,m$,
which is referred to as the positivity assumption.
The average treatment effect (ATE) is defined as $\mathbb{E}[Y^{(a=1)}-Y^{(a=0)}]$. This parameter captures the difference in the potential outcome mean if all units received treatment $A=1$ and if all units received treatment $A=0$ (e.g., placebo treatment). In mediation analysis, the focus is on the fact that this causal effect is partly mediated through the variable $M$, and it is of interest to quantify the direct and indirect portions of the causal effect by partitioning the ATE as follows (Robins and Greenland, , 1992; Pearl, , 2001).
$\mathbb{E}\left[Y^{(1)}-Y^{(0)}\right]=\mathbb{E}\left[Y^{\left(1,M^{(1)}\right)}-Y^{\left(0,M^{(0)}\right)}\right]=\mathbb{E}\left[Y^{\left(1,M^{(1)}\right)}-Y^{\left(1,M^{(0)}\right)}\right]+\mathbb{E}\left[Y^{\left(1,M^{(0)}\right)}-Y^{\left(0,M^{(0)}\right)}\right].$ (2.1)
The first and the second terms in the last expression are called the total indirect effect and the pure direct effect, respectively by Robins and Greenland, (1992), and are called the natural indirect effect (NIE) and the natural direct effect (NDE) of the treatment on the outcome, respectively by Pearl, (2001). We will use the latter terminology in this work. NIE captures the change in the expectation of the response variable in a hypothetical scenario where the value of the treatment variable is fixed at $A=1$, while the mediator behaves as if the treatment had been changed from value 0 to 1. NDE captures the change in the expectation of the outcome in a hypothetical scenario where the value of the treatment variable is changed from value 0 to 1, while the mediator behaves as if the treatment is fixed at 0. Note that if one can identify and estimate the parameter $\mathbb{E}[Y^{(a^{\prime},M^{(a)})}]$ for $a,a^{\prime}\in\{0,1\}$, then based on Equation (2.1), NIE, NDE and ATE can all be estimated. Henceforth, we focus on the parameter $\mathbb{E}[Y^{(1,M^{(0)})}]$ in this paper.
We operate under the assumption of sequential exchangeability, which implies there are no unobserved confounders for the treatment-outcome, mediator-outcome, and treatment-mediator relationships (Imai et al., , 2010), formalized as follows.
Assumption 2.1 (Sequential exchangeability).
Let $X_{1}\perp\!\!\!\!\perp X_{2}\mid X_{3}$ indicate that the random variables $X_{1}$ and $X_{2}$ are conditionally independent given the random variable $X_{3}$.
  1. 1.
    $Y^{(a,m)}\perp\!\!\!\!\perp\{A,M\}\mid X$, for all $a,m$,
  2. 2.
    $M^{(a)}\perp\!\!\!\!\perp A\mid X$, for all $a$,
  3. 3.
    $Y^{(a,m)}\perp\!\!\!\!\perp M^{(a^{\prime})}\mid X$, for all $a,a^{\prime},m$.
As established in (Imai et al., , 2010), under Assumption 2.1, the parameter $\mathbb{E}[Y^{(1,M^{(0)})}]$ can be identified as
[…]
where we use $f$ to represent densities over variables. The right hand side is known as the mediation functional in the literature, and will be our parameter of interest in this work.
The estimation of the mediation effect is a well-studied problem in statistics in both parametric and non-parametric setups. Yet, in the presence of UHD variables with both $p$ and $q$ larger than the sample size, the preference often leans toward simple linear models instead of intricate non-linear models due to the developed theory as well as the interpretability of the model. Hence, we focus on the following data generating mechanisms for the outcome and mediator variables:
$Y$ $=(1-A)\left(\alpha_{0}+X^{\top}\beta_{0}+M^{\top}\gamma_{0}\right)+A\left(\alpha_{1}+X^{\top}\beta_{1}+M^{\top}\gamma_{1}\right)+\epsilon,$ (2.3)
$M$ $=(1-A)\left(\delta_{0}+\mathbf{B}_{0}X\right)+A\left(\delta_{1}+\mathbf{B}_{1}X\right)+U,$
where $\epsilon$ and $U$ are centered random shocks and the matrices $\mathbf{B}_{1},\mathbf{B}_{0}\in\mathbb{R}^{q\times p}$, and vectors $\beta_{1},\beta_{0}\in\mathbb{R}^{p}$, $\gamma_{1},\gamma_{0}\in\mathbb{R}^{q}$ contain unknown regression coefficients that encode the relationships among the exposures, mediators and outcome. Note that this data generating process (DGP) is flexible in the sense that
  1. 1.
    it allows for UHD mediators $M$ and UHD pre-treatment covariates $X$,
  2. 2.
    it does not require any particular modeling assumptions or restrictions on the propensity score, which is that conditional distribution of the treatment variable $A$ given the covariates $X$, and
  3. 3.
    it allows for $A-X$ and $A-M$ interactions (although it requires no $X-M$ interaction).
This DGP implies the following.
$\mu_{1}(X,M):=\mathbb{E}[Y\mid X,M,A=1]=X^{\top}\beta_{1}+M^{\top}\gamma_{1},$ (2.4)
$\mu_{10}(X):=\mathbb{E}[\mu_{1}(X,M)\mid X,A=0]=X^{\top}\beta_{1}+X^{\top}\mathbf{B}_{0}^{\top}\gamma_{1},$ (2.5)
for all $X\in\mathbb{R}^{p},M\in\mathbb{R}^{q}$, where we absorbed the intercepts in the means. Therefore, the parameter of interest in Equation (2.2) reduced to
$\mathbb{E}[\mu_{10}(X)]=\mathbb{E}[X]^{\top}(\beta_{1}+\mathbf{B}_{0}^{\top}\gamma_{1}).$
We denote the number of treated units by $n_{t}$, and the number of control units by $n_{c}$. We use the notations $\mathbf{X}_{c}$, $\mathbf{X}_{t}$ and $\mathbf{M}_{c}$, $\mathbf{M}_{t}$ to refer to the feature matrices that solely correspond to control or treated units, respectively, with different rows corresponding to the observations of different individuals. Similar to Athey et al., (2018), we focus on the sample version of the parameter of interest, which for the mediation functional, will be
$\theta_{0}=\frac{1}{n}\sum_{i=1}^{n}\mu_{10}(X_{i})=\bar{X}^{\top}(\beta_{1}+\mathbf{B}_{0}^{\top}\gamma_{1}),$ (2.6)
where $\bar{X}=\frac{1}{n}\sum_{i=1}^{n}X_{i}$. In the next section, we will describe our proposed methodology for estimating $\theta_{0}$.
3 Debiased Estimator for the Mediation Functional
In this section, we present our methodology for estimating the mediation functional. A naive approach for estimating our parameter of interest $\theta_{0}=\bar{X}^{\top}\left(\beta_{1}+\mathbf{B}_{0}^{\top}\gamma_{1}\right)$ is by simply using the plug-in method: in the first step, we estimate $(\beta_{1},\gamma_{1})$ by regressing $Y$ on $(M,X)$ using the treatment observations, and then, we estimate $\mathbf{B}_{0}$ by regressing $M$ on $X$ (coordinate-wise) using control observations. Thus, our estimator becomes $\hat{\theta}^{\rm plug-in}=\bar{X}^{\top}(\hat{\beta}_{1}+\hat{\mathbf{B}}_{0}^{\top}\hat{\gamma}_{1})$. As the setting under consideration is UHD, to achieve consistent estimation, we can employ a penalized regression method, such as Lasso. However, this will introduce bias into the estimation process through $(\hat{\beta}_{1},\hat{\gamma}_{1},\hat{\mathbf{B}}_{0})$. Hence, our goal is to devise a debiasing procedure for this estimator so that the refined estimator is $\sqrt{n}$-consistent and asymptotically normal.
To motivate our debiasing procedure, we first consider a simpler scenario, where we assume $\mathbf{B}_{0}$ is known. In that case, we can simply debias $(\hat{\beta}_{1},\hat{\gamma}_{1})$ in the direction of $(\bar{X},\mathbf{B}_{0}\bar{X})$ to obtain a debiased estimator. Using tools from the literature on the standard high-dimensional debiased Lasso estimator, it is easy to prove that the resulting estimator is $\sqrt{n}$-consistent and asymptotically normal. However, this debiasing procedure is not fruitful when $\mathbf{B}_{0}$ is unknown. One might substitute $\mathbf{B}_{0}$ by $\hat{\mathbf{B}}_{0}$ (obtained through penalized regression of $M$ on $X$ coordinate-wise), but as mentioned previously, it is likely to be a biased estimator of $\mathbf{B}_{0}$ due to penalization, and consequently, our direction to debias $(\hat{\beta}_{1},\hat{\gamma}_{1})$ itself becomes biased. Hence, another debiasing step is needed.
[…]
Figure 1: The proposed estimation procedure.
We now present the details of our methodology, which is summarized in Algorithm 1 and in Figure 1. Our method essentially comprises five key steps. In the first step, we obtain $\hat{\beta}_{1},\hat{\gamma}_{1}$ by regressing $Y$ on $(X,M)$ using the treatment observations. As we are in a UHD regime, we use the $\ell_{1}$ penalty for consistent estimation. We also regress $M$ on $X$ using control observations, along with $\ell_{1}$ penalty, to obtain $\hat{\mathbf{B}}_{0}$. As $M$ is multivariate, we regress each coordinate of $M$ on $X$ in parallel to estimate the rows of $\mathbf{B}_{0}$.
Algorithm 1 Estimation procedure
Step 1: Split the data into two parts, say $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$. Use $\mathcal{D}_{1}$ to run $q$ parallel regressions of $(\mathbf{M}_{c})_{.,j},j=1,\ldots,q$ on $\mathbf{X}_{c}$, to obtain $\hat{\mathbf{B}}_{0_{1,.}},\ldots,\hat{\mathbf{B}}_{0_{q,.}}$ where each $\hat{\mathbf{B}}_{0_{i,.}}\in{\mathbb{R}}^{p}$. Concatenate them to obtain the following estimator for $\mathbf{B}_{0}$
$\hat{\mathbf{B}}_{0}=[\hat{\mathbf{B}}_{0_{1,.}}\quad\hat{\mathbf{B}}_{0_{2,.}}\quad\cdots\quad\hat{\mathbf{B}}_{0_{q,.}}]^{\top}.$
Use $\mathcal{D}_{2}$ to Estimate ${\phi}_{1}=[{\beta}_{1}^{\top}\quad{\gamma}_{1}^{\top}]^{\top}$ by regressing $Y_{t}$ on $\mathbf{W}_{t}=[\mathbf{X}_{t}\quad\mathbf{M}_{t}]$ along with a $\ell_{1}$ penalty
$\hat{\phi}_{1}=[{\beta}_{1}^{\top}\quad{\gamma}_{1}^{\top}]^{\top}=\arg\min_{\tilde{\phi}_{1}}\left\{\sum_{\{i:A_{i}=1\}}\left(Y_{i}-\mathbf{W}_{i,.}\cdot\tilde{\phi}_{1}\right)^{2}+\lambda_{1}\|\tilde{\phi}_{1}\|_{1}\right\}.$
Step 2: Compute bias correction weights $\tau_{1}$ with contrast $a=\left[\bar{X}^{\top}\quad(\hat{\mathbf{B}}_{0}\bar{X})^{\top}\right]^{\top}$ and tuning parameter $K_{1}$ using $\mathcal{D}_{2}$,
$\tau_{1}=\operatorname{argmin}_{\tilde{\tau}_{1}}\left\{\left\lVert\tilde{\tau}_{1}\right\rVert_{2}^{2}\quad\text{subject to}\quad\left\lVert a-\mathbf{W}_{t}^{\top}\tilde{\tau}_{1}\right\rVert_{\infty}\leqslant K_{1}\sqrt{\frac{\log(p+q)}{n_{t}}},\|\tilde{\tau}_{1}\|_{\infty}\leq n_{t}^{-2/3}\right\},$ (3.1)
and set
$\hat{\theta}_{0,1}=a^{\top}\hat{\phi}_{1}+\sum_{\{i:A_{i}=1\}}\tau_{1,i}\left(Y_{i}-\mathbf{W}_{i,.}\cdot\hat{\phi}_{1}\right).$
Step 3: Estimate $b=\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}$ by regressing $M_{i}^{\top}\hat{\gamma}_{1}$ on $X_{i}$ (with $A_{i}=0$) along with a $\ell_{1}$ penalty using $\mathcal{D}_{2}$:
$\hat{b}=\arg\min_{\tilde{b}}\left\{\sum_{\{i:A_{i}=0\}}\left(M_{i}^{\top}\hat{\gamma}_{1}-X_{i}^{\top}\tilde{b}\right)^{2}+\lambda_{2}\|\tilde{b}\|_{1}\right\}.$
Step 4: Compute bias correction weight $\tau_{2}$ with contrast $\bar{X}$ and tuning parameter $K_{2}$ using $\mathcal{D}_{2}$,
$\tau_{2}=\operatorname{argmin}_{\tilde{\tau}_{2}}\left\{\|\tilde{\tau}_{2}\|_{2}^{2}\quad\text{ subject to }\quad\|\bar{X}-\mathbf{X}_{c}^{\top}\tilde{\tau}_{2}\|_{\infty}\leqslant K_{2}\sqrt{\frac{\log(p)}{n_{c}}},\|\tilde{\tau}_{2}\|_{\infty}\leq n_{c}^{-2/3}\right\},$ (3.2)
and set
$\hat{\theta}_{0,2}=\bar{X}^{\top}\hat{b}+\sum_{\{i:A_{i}=0\}}\tau_{2,i}\left(M_{i}^{\top}\hat{\gamma}_{1}-X_{i}^{\top}\hat{b}\right).$
Step 5: Return the final estimator $\hat{\theta}_{0}=\hat{\theta}_{0,1}+\hat{\theta}_{0,2}-(\hat{\mathbf{B}}_{0}\bar{X})^{\top}\hat{\gamma}_{1}$.
In the second step, we debias $\hat{\phi}_{1}:=(\hat{\beta}_{1},\hat{\gamma}_{1})$ in the direction of $a:=(\bar{X}\quad\hat{\mathbf{B}}_{0}\bar{X})$ to obtain an intermediate estimator $\hat{\theta}_{0,1}$. For notational simplicity, we define $\mathbf{W}=(\mathbf{X}\quad\mathbf{M})$ and $\mathbf{W}_{t}=(\mathbf{X}_{t}\quad\mathbf{M}_{t})$. More specifically, we have:
$\hat{\theta}_{0,1}=\bar{X}^{\top}\hat{\beta}_{1}+(\hat{\mathbf{B}}_{0}\bar{X})^{\top}\hat{\gamma}_{1}+\sum_{\{i:A_{i}=1\}}\tau_{1,i}\left(Y_{i}-\mathbf{W}_{i,.}\cdot\hat{\phi}_{1}\right)\ \equiv\ \hat{\theta}^{\rm plug-in}+\sum_{\{i:A_{i}=1\}}\tau_{1,i}\left(Y_{i}-\mathbf{W}_{i,.}\cdot\hat{\phi}_{1}\right)\,,$
where the additional summand is the correction term obtained by solving a suitable optimization problem (see Equation (3.1) for details) that aims to reduce the bias of the penalized estimator. As mentioned earlier, this estimator is not completely free of bias due to the estimated $\hat{\mathbf{B}}_{0}$.
To remove this bias, in the third step, we first regress $M^{\top}\hat{\gamma}_{1}$ on $X$ along with an $\ell_{1}$ penalty (as $X$ is a UHD vector). As $\mathbb{E}[M\mid X,A=0]=\mathbf{B}_{0}X$, this penalized regression effectively estimates $\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}$. In the fourth step, we debias this estimator along the direction of $\bar{X}$, i.e., we obtain
$\hat{\theta}_{0,2}=\bar{X}\cdot\widehat{\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}}+\sum_{\{i:A_{i}=0\}}\tau_{2,i}\left(M_{i}^{\top}\hat{\gamma}_{1}-X_{i}^{\top}\hat{b}\right),$
where $b:=\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}$, and $\hat{b}=\widehat{\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}}$, and the additional summand is the correction term obtained by solving a suitable optimization problem (see Equation (3.2) for details) that aims to reduce the bias of the penalized estimator. In the final step, we obtain our final debiased estimator $\hat{\theta}_{0}$ as
$\hat{\theta}_{0}=\hat{\theta}_{0,1}+\hat{\theta}_{0,2}-(\hat{\mathbf{B}}_{0}\bar{X})^{\top}\hat{\gamma}_{1}\,.$ (3.3)
As will be seen in Proposition 4.1, the interplay between $\widehat{\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}}$ and $\hat{\mathbf{B}}_{0}^{\top}\hat{\gamma}_{1}$ leads to a second order bias for $\hat{\theta}_{0}$. In the next two sections, we delve into the analysis of our proposed estimator. We provide results regarding the structure of the bias of the estimator in Section 4 and demonstrate that the estimator is $\sqrt{n}$-consistent and asymptotically normal in Section 5.
4 Bias Analysis and Connection to Fixed-Dimensional Setting
4.1 Bias Analysis
In this section, we present the bias structure of the debiased estimator $\hat{\theta}_{0}$ and show that the bias is asymptotically negligible, i.e., it decays at a faster rate than $n^{-1/2}$. We would like to point out that debiasing a high-dimensional parameter does not make it unbiased, i.e., it does not remove the bias completely, rather reduces the bias enough to achieve $\sqrt{n}$-consistency (e.g., see (Zhang and Zhang, , 2014; Van de Geer et al., , 2014; Javanmard and Montanari, , 2014)). To analyze the remaining bias of $\hat{\theta}_{0}$, first, recall the definitions $b:=\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}$, $\hat{b}:=\widehat{\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}}$, $\hat{\phi}_{1}:=(\hat{\beta}_{1}^{\top},\hat{\gamma}_{1}^{\top})^{\top}$, and the contrast $a:=(\bar{X}^{\top}\quad(\hat{\mathbf{B}}_{0}\bar{X})^{\top})^{\top}$ from the previous section. The following proposition establishes an upper bound on the bias of $\hat{\theta}_{0}$.
Proposition 4.1.
The estimation error of the debiased estimator $\hat{\theta}_{0}$ can be decomposed as
$\hat{\theta}_{0}-\theta_{0}=\Delta_{n}+V_{n},$
where,
$V_{n}=\sum_{\{i:A_{i}=1\}}\tau_{1,i}\epsilon_{i}+\sum_{\{i:A_{i}=0\}}\tau_{2,i}U_{i}^{\top}\hat{\gamma}_{1}$
is a centered random variable, and $\Delta_{n}$ is the remaining bias, which can be bounded as
$|\Delta_{n}|\leq\|\mathbf{B}_{0}\bar{X}-\hat{\mathbf{B}}_{0}\bar{X}\|_{\infty}\|\hat{\gamma}_{1}-\gamma_{1}\|_{1}+\|a-\mathbf{W}_{t}^{\top}\tau_{1}\|_{\infty}\|\hat{\phi}_{1}-\phi_{1}\|_{1}+\|\bar{X}-\mathbf{X}_{c}^{\top}\tau_{2}\|_{\infty}\|\hat{b}-b\|_{1}\,.$ (4.1)
See Appendix A.1 for a proof. Observe that each of the three terms in the upper bound of $|\Delta_{n}|$ has a product form, i.e., the product of an estimation and/or an approximation error. More specifically, the first summand is the product of the estimation error of $\mathbf{B}_{0}$ (along the direction of $\bar{X}$) and $\gamma_{1}$, the second summand is the product of the estimation error of $\phi_{1}$ and the approximation error of $a$ by $\mathbf{W}_{t}^{\top}\tau_{1}$, and the last summand is the product of the estimation error of $b$ and the approximation error of $\bar{X}$ by $\mathbf{X}_{c}^{\top}\tau_{2}$. Whilst each of these individual estimation/approximation errors can be significantly larger than $n^{-1/2}$ in order, the product form is what makes each summand asymptotically negligible. In Section 5, we provide sufficient conditions on the sparsity of parameters and the distribution of the error terms, under which $|\Delta_{n}|=o_{p}(n^{-1/2})$, and $\hat{\theta}_{0}$ is $\sqrt{n}$-consistent and asymptotically normal.
4.2 Connection to Fixed-Dimensional Setting
In a fixed dimensional setting (i.e., when the number of covariates as well as potential mediators are fixed), Tchetgen Tchetgen and Shpitser, (2012) provided a debiased estimator of the mediation functional which was constructed based on the influence function of the parameter of interest. Our proposed estimator can be considered as the counterpart of their approach in the UHD setting. The debiased estimator, as proposed in (Tchetgen Tchetgen and Shpitser, , 2012), can be written as
$\textstyle\hat{\theta}^{IF}$ $=\mathbb{E}_{n}\bigg{[}\hat{\mu}_{10}(X)+\frac{A}{\hat{p}(A=1\mid X)}\frac{\hat{p}(M\mid A=0,X)}{\hat{p}(M\mid A=1,X)}\Big{(}Y-\hat{\mu}_{1}(X,M)\Big{)}+\frac{1-A}{\hat{p}(A=0\mid X)}\Big{(}\hat{\mu}_{1}(X,M)-\hat{\mu}_{10}(X)\Big{)}\bigg{]}\,,$ (4.2)
[…]
We now show that our estimator $\hat{\theta}_{0}$, obtained via Algorithm 1, is structurally similar to $\hat{\theta}^{IF}$ when the underlying data generating process is linear. Consider the following specific nuisance function choices $\hat{\mu}_{1}(X,M)=X^{\top}\hat{\beta}_{1}+M^{\top}\hat{\gamma}_{1}$, and $\hat{\mu}_{10}(X)=X^{\top}(\hat{\beta}_{1}+\widehat{\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}})$. Then, estimator $\hat{\theta}^{IF}$ will be of the form of
$\hat{\theta}^{IF}$ $=\mathbb{E}_{n}\bigg{[}X^{\top}\left(\hat{\beta}_{1}+\widehat{\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}}\right)+\frac{A}{\hat{\pi}(X)}\frac{\hat{p}(M\mid A=0,X)}{\hat{p}(M\mid A=1,X)}\left(Y-X^{\top}\hat{\beta}_{1}-M^{\top}\hat{\gamma}_{1}\right)$
$\quad+\frac{1-A}{1-\hat{\pi}(X)}\left(X^{\top}\hat{\beta}_{1}+M^{\top}\hat{\gamma}_{1}-X^{\top}\left(\hat{\beta}_{1}+\widehat{\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}}\right)\right)\bigg{]}$
$=\mathbb{E}_{n}\left[X^{\top}\left(\hat{\beta}_{1}+\widehat{\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}}\right)+\omega_{1}A\left(Y-X^{\top}\hat{\beta}_{1}-M^{\top}\hat{\gamma}_{1}\right)+\omega_{2}(1-A)\left(M^{\top}\hat{\gamma}_{1}-X^{\top}\widehat{\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}}\right)\right]$
$=\bar{X}^{\top}\left(\hat{\beta}_{1}+\widehat{\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}}\right)+\frac{1}{n}\sum_{\{i:A_{i}=1\}}\omega_{1,i}\left(Y_{i}-X_{i}^{\top}\hat{\beta}_{1}-M_{i}^{\top}\hat{\gamma}_{1}\right)+\frac{1}{n}\sum_{\{i:A_{i}=0\}}\omega_{2,i}\left(M_{i}^{\top}\hat{\gamma}_{1}-X_{i}^{\top}\widehat{\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}}\right)\,,$ (4.3)
where
$\hat{\pi}(X_{i})=\hat{p}(A=1\mid X_{i}),\ \ \omega_{1,i}=\frac{1}{\hat{\pi}(X_{i})}\frac{\hat{p}(M_{i}\mid A_{i}=0,X_{i})}{\hat{p}(M_{i}\mid A_{i}=1,X_{i})},\quad\omega_{2,i}=\frac{1}{1-\hat{\pi}(X_{i})}.$
As it can be seen, the estimator $\hat{\theta}^{IF}$ contains a regression-based initial estimator, along with two weighted terms for bias correction. This is the same structure as our proposed estimator:
$\hat{\theta}_{0}=\bar{X}^{\top}(\hat{\beta}_{1}+\widehat{\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}})+\sum_{\{i:A_{i}=1\}}\tau_{1,i}\left(Y_{i}-X_{i}^{\top}\hat{\beta}_{1}-M_{i}^{\top}\hat{\gamma}_{1}\right)+\sum_{\{i:A_{i}=0\}}\tau_{2,i}\left(M_{i}^{\top}\hat{\gamma}_{1}-X_{i}^{\top}\widehat{\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}}\right)\,$ (4.4)
It is evident from Equation (4.2) and Equation (4.4) that our estimator possesses the same form as $\hat{\theta}^{IF}$ with $\tau_{1}$ and $\tau_{2}$ playing the role of weights $\omega_{1}/n$ and $\omega_{2}/n$, respectively.
We next consider the following reparametrization of the estimator of the mediation functional in Equation (4.2), as recently studied in (Ghassami, , 2024):
$\hat{\theta}^{mr}=\mathbb{E}_{n}\left[\hat{\mu}_{10}(X)+\frac{(1-A)}{\hat{\varpi}_{1}(X)}\left\{\hat{\mu}_{1}(M,X)-\hat{\mu}_{10}(X)\right\}+\frac{A\hat{\varpi}_{2}(M,X)}{1-\hat{\varpi}_{1}(X)}\left\{Y-\hat{\mu}_{1}(M,X)\right\}\right],$ (4.5)
where,
$\varpi_{1}(X)$ $=p(A=0\mid X),\quad\varpi_{2}(M,X)=\frac{p(M\mid A=0,X)}{p(M\mid A=1,X)}.$
It was shown in (Ghassami, , 2024) that under this reparametrization in the fixed-dimensional setting, the bias has the following form.
$\mathbb{E}\left[\hat{\theta}^{mr}\right]-\theta_{0}=$ $\mathbb{E}\left[\left\{\frac{(1-A)}{\hat{\varpi}_{1}(X)}-1\right\}\left\{\mu_{10}(X)-\hat{\mu}_{10}(X)\right\}\right]$
$\qquad+\mathbb{E}\left[\left\{\frac{A\hat{\varpi}_{2}(M,X)}{1-\hat{\varpi}_{1}(X)}-\frac{(1-A)}{\hat{\varpi}_{1}(X)}\right\}\left\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\right\}\right].$ (4.6)
In the fixed-dimension, by the Hölder’s inequality and plugging in $\hat{\mu}_{1}(X,M)=X^{\top}\hat{\beta}_{1}+M^{\top}\hat{\gamma}_{1}$, and $\hat{\mu}_{10}(X)=X^{\top}(\hat{\beta}_{1}+\widehat{\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}})$, Equation (4.6) can be upper bounded by
$\left|\mathbb{E}[\hat{\theta}^{mr}]-\theta_{0}\right|$
$\leq\mathbb{E}\left[\left\{\left\|X^{\top}\left(\frac{A\hat{\varpi}_{2}(M,X)}{1-\hat{\varpi}_{1}(X)}-\frac{1-A}{\hat{\varpi}_{1}(X)}\right)\right\|_{\infty}+\left\|X^{\top}\left(\frac{1-A}{\hat{\varpi}_{1}(X)}-1\right)\right\|_{\infty}\right\}\|\hat{\beta}_{1}-\beta_{1}\|_{1}\right]$
$\quad+\mathbb{E}\left[\left\|M^{\top}\left(\frac{A\hat{\varpi}_{2}(M,X)}{1-\hat{\varpi}_{1}(X)}-\frac{(1-A)}{\hat{\varpi}_{1}(X)}\right)\right\|_{\infty}\left\|\gamma_{1}-\hat{\gamma}_{1}\right\|_{1}\right]$
$\quad+\mathbb{E}\left[\left\|X^{\top}\left(\frac{1-A}{\hat{\varpi}_{1}(X)}-1\right)\right\|_{\infty}\left\|\mathbf{B}_{0}^{\top}\gamma_{1}-\widehat{\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}}\right\|_{1}\right].$ (4.7)
Now, coming back to our estimator $\hat{\theta}_{0}$, the remaining bias $\Delta_{n}$, as shown in Proposition 4.1, is bounded by:
$|\Delta_{n}|$ $\leq\|a-\mathbf{W}^{\top}\tau_{1}\|_{\infty}\|\hat{\phi}_{1}-\phi_{1}\|_{1}+\|(\hat{\mathbf{B}}_{0}-\mathbf{B}_{0})\bar{X}\|_{\infty}\|\hat{\gamma}_{1}-\gamma_{1}\|_{1}+\|\bar{X}-\mathbf{X}_{c}^{\top}\tau_{2}\|_{\infty}\|\hat{b}-b\|_{1}$
$=\|a-\mathbf{W}^{\top}\tau_{1}\|_{\infty}\|\hat{\beta}_{1}-\beta_{1}\|_{1}$
$\quad+\left(\|a-\mathbf{W}^{\top}\tau_{1}\|_{\infty}+\|(\hat{\mathbf{B}}_{0}-\mathbf{B}_{0})\bar{X}\|_{\infty}\right)\|\hat{\gamma}_{1}-\gamma_{1}\|_{1}$
$\quad+\|\bar{X}-\mathbf{X}_{c}^{\top}\tau_{2}\|_{\infty}\|\widehat{\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}}-\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}\|_{1}.$
It is apparent that the bias structures of $\hat{\theta}^{mr}$ and $\hat{\theta}_{0}$ are similar, as both of them are the sum of three terms, and each of them is the product of two errors. We summarize the similarity in the following table.
$\begin{array}[]{c c c}\text{High-dimensional setting}&&\text{Fixed-dimensionalsetting}\\ \hline\cr\|a-\mathbf{W}^{\top}\tau_{1}\|_{\infty}\|\hat{\beta}_{1}-\beta_{1}\|_{1}&\longleftrightarrow&\begin{matrix}\big{(}\|X^{\top}\left(\frac{A\hat{\varpi}_{2}(M,X)}{1-\hat{\varpi}_{1}(X)}-\frac{1-A}{\hat{\varpi}_{1}(X)}\right)\|_{\infty}\\ ~{}~{}~{}~{}~{}~{}+\|X^{\top}\left(\frac{1-A}{\hat{\varpi}_{1}(X)}-1\right)\|_{\infty}\big{)}\|\beta_{1}-\hat{\beta}_{1}\|_{1}\end{matrix}\\ \left(\|a-\mathbf{W}^{\top}\tau_{1}\|_{\infty}+\|(\hat{\mathbf{B}}_{0}-\mathbf{B}_{0})\bar{X}\|_{\infty}\right)\|\hat{\gamma}_{1}-\gamma_{1}\|_{1}&\longleftrightarrow&\|M^{\top}\left(\frac{A\hat{\varpi}_{2}(M,X)}{1-\hat{\varpi}_{1}(X)}-\frac{(1-A)}{\hat{\varpi}_{1}(X)}\right)\|_{\infty}\left\|\gamma_{1}-\hat{\gamma}_{1}\right\|_{1}\\ \|\bar{X}-\mathbf{X}_{c}^{\top}\tau_{2}\|_{\infty}\|\widehat{\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}}-\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}\|_{1}&\longleftrightarrow&\|X^{\top}\left(\frac{1-A}{\hat{\varpi}_{1}(X)}-1\right)\|_{\infty}\|\mathbf{B}_{0}^{\top}\gamma_{1}-\widehat{\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}}\|_{1}\\ \end{array}$
Relation to balancing: We also draw a connection between our debiasing technique and the balancing techniques in the literature of causal inference. In the influence function-based estimator for ATE, the correction term (the term added to the initial estimator to reduce bias) contains the propensity score as a nuisance function. Imai and Ratkovic, (2014) showed that the propensity score satisfies a certain balancing property, based on which they proposed a balancing estimator for the propensity score. Zubizarreta, (2015) proposed an optimization-based approach for achieving the said balance. In the high-dimensional setting, Athey et al., (2018) illustrated that the standard debiasing technique of penalized regression-based estimator (e.g., the method proposed earlier in (Javanmard and Montanari, , 2014; Van de Geer et al., , 2014; Zhang and Zhang, , 2014)) is inherently related to the balancing approach of Zubizarreta, (2015). However, none of these approaches is directly applicable in the presence of mediators for assessing the mediation functional. Recently, Ghassami, (2024) presented a counterpart of the balancing property and balancing estimator of Imai and Ratkovic, (2014) for mediation functional in a fixed-dimensional setting. A version of the balancing property in that work can be achieved by minimizing $\left\|\frac{1-A}{\hat{\varpi}_{1}(X)}-1\right\|_{\infty}$ and $\left\|\frac{A\hat{\varpi}_{2}(M,X)}{1-\hat{\varpi}_{1}(X)}-\frac{(1-A)}{\hat{\varpi}_{1}(X)}\right\|_{\infty}$. As described above, the counter part of these two quantities in our setting are $\|\bar{X}-\mathbf{X}_{c}^{\top}\tau_{2}\|_{\infty}$ and $\|a-\mathbf{W}_{t}^{\top}\tau_{1}\|_{\infty}$, respectively. Hence, the two optimizations (3.1) and (3.2) performed in our proposed estimation strategy can be viewed as adaptations of the balancing of Ghassami, (2024) in the UHD setting.
5 Theoretical Analysis
5.1 Assumptions
We define the sub-Gaussian and sub-Exponential norms of a random variable $X$, which will be needed in the analysis of our proposed approach.
Definition 5.1.
The sub-Gaussian norm of a random variable $X$, denoted by $\|X\|_{\psi_{2}}$, is defined as
$\|X\|_{\psi_{2}}=\sup_{q\geq 1}q^{-1}(\mathbb{E}|X|^{q})^{1/q}.$
For a random vector $X\in\mathbb{R}^{n}$, its sub-Gaussian norm is defined as
$\|X\|_{\psi_{2}}=\sup_{x\in S^{n-1}}\|\langle X,x\rangle\|_{\psi_{2}},$
where $S^{n-1}$ denotes the unit sphere in $\mathbb{R}^{n}$. For any $t>0$, the moment generating generating function of a sub-Gaussian random variable satisfies
$\mathbb{E}[{\sf exp}(t(X_{i}-\mathbb{E}[X_{i}]))]\leq{\sf exp}(\varrho^{2}t^{2}/2)$
where $\varrho$ is the sub-Gaussian constant.
Definition 5.2.
The sub-Exponential norm of a random variable $X$, denoted by $\|X\|_{\psi_{1}}$, is defined as
$\|X\|_{\psi_{1}}=\sup_{q\geq 1}q^{-1/2}(\mathbb{E}|X|^{q})^{1/q}.$
For a random vector $X\in\mathbb{R}^{n}$, its sub-Gaussian norm is defined as
$\|X\|_{\psi_{1}}=\sup_{x\in S^{n-1}}\|\langle X,x\rangle\|_{\psi_{1}},$
where $S^{n-1}$ denotes the unit sphere in $\mathbb{R}^{n}$.
Definition 5.3 (Restricted eigenvalue condition and compatibility constant).
Let $\mathbf{X}$ be an $n\times p$ design matrix with i.i.d. rows $\{X_{1},\dots,X_{n}\}$ and let $\operatorname{Var}(X_{i})=\Sigma_{X}$. Given the sparsity level $k_{0}$, we say that $\Sigma_{X}$ satisfies $\{k_{0},\eta_{0},L_{0}\}-$restricted eigenvalue condition, with some $\eta>0$, if $\nu^{\top}\Sigma_{X}\nu\geq\eta_{0}\|\nu\|_{2}^{2}$ for all $\nu\in\mathcal{C}_{k_{0}}(L_{0})$, where $\mathcal{C}_{k_{0}}(L_{0})$ is defined as, for $1\leq k_{0}\leq p$ and $L_{0}\geq 1$,
$\mathcal{C}_{k_{0}}(L_{0})=\left\{\nu\in\mathbb{R}^{p}:\|\nu\|_{1}\leq L_{0}\sum_{j=1}^{k_{0}}\left|\nu_{i_{j}}\right|\text{ for some }1\leq i_{1}<\ldots<i_{j}\leq p\right\}$ (5.1)
The corresponding compatibility constant is defined as:
$\varphi_{X}^{2}(\Sigma_{X},\eta_{0},L_{0}):=\min_{\nu\in\mathbb{R}^{p}}\left\{\frac{\nu^{\top}\Sigma_{X}\nu}{\eta_{0}\|\nu\|_{1}^{2}}:\nu\in\mathbb{R}^{p},\nu\in\mathcal{C}_{k_{0}}(L_{0})\right\}$ (5.2)
We now study the theoretical properties of the estimator proposed in Equation (3.3). We begin by stating the required assumptions for the theoretical analysis.
Assumption 5.4 (Sparsity).
We assume that $\beta_{1}$ is $k_{1}$-sparse, i.e., $\|\beta_{1}\|_{0}\leqslant k_{1}$, $\gamma_{1}$ is $k_{2}$-sparse, i.e., $\|\gamma_{1}\|_{0}\leqslant k_{2}$ (which implies $\phi_{1}$ is $k_{1}+k_{2}=k$ sparse), and $\|\hat{\phi}_{1}\|_{0}\leqslant 2(k_{1}+k_{2})$ with high probability [^1] In Step 1 of Algorithm 1, when performing lasso regression of $Y$ on $(X,M)$ on the treated units, we assume that $\hat{\beta}_{1}$ and $\hat{\gamma}_{1}$ are $2k_{1}$-sparse and $2k_{2}$-sparse, respectively, with probability approaching to 1. This implies that $\hat{\phi}_{1}$ is $2(k_{1}+k_{2})$-sparse with high probability. Since the lasso leads to sparsity with high probability, this assumption removes the need for a stronger condition on bounding the $\ell_{1}$-risk of $b-\hat{b}$, which would otherwise require $\mathbf{B}_{0}$ to exhibit column sparsity. . $\mathbf{B}_{0}$ is $s$-row sparse (each row of $\mathbf{B}_{0}$ has at most $s$ active element, i.e., $\|\mathbf{B}_{0,i}\|_{0}\leqslant s$). Furthermore, $(k\vee s)\log(p+q)/\sqrt{n}=o(1)$ and $sk\log(p\vee q)/\sqrt{n_{c}\wedge n_{t}}=o(1)$.
The assumption $(k\vee s)\log(p+q)/\sqrt{n}=o(1)$ typically arises in debiased lasso literature (Javanmard and Montanari, , 2014; Van de Geer et al., , 2014; Zhang and Zhang, , 2014) to establish the $\sqrt{n}$-consistency of the debiased estimator. Cai and Guo, (2017) proved that this assumption is necessary unless we have more information on the design. The last assumption, i.e., $sk\log(p\vee q)/\sqrt{n_{c}\wedge n_{t}}=o(1)$, is required for the high-dimensional matrix over matrix regression (Step $1$ in our algorithm 1) and further bias correction. Whether this is necessary, we leave this as a future work.
Assumption 5.5 (Sub-Gaussian condition).
Assume that the noise is homoskedastic: $\operatorname{Var}[\epsilon_{i}^{(1)}|\mathbf{W}_{t,i}]=\sigma_{1}^{2}$ for every $i=1,\ldots,n_{t}$, and that the noise $\epsilon_{i}^{(1)}:=Y_{i}^{(1)}-\mathbb{E}[Y_{i}^{(1)}|\mathbf{W}_{t,i}]$ is i.i.d uniformly sub-Gaussian with mean 0 a finite parameter $v_{1}^{2}S_{1}>0$. Similarly, suppose that we have homoskedastic noise: $\operatorname{Var}[U_{i*}^{(0)}\mid X_{i}]=\sigma^{2}_{2}$ for all $i=1,\ldots,n_{c}$, and also the response noise $U_{i*}^{(0)}$ is i.i.d uniformly sub-Gaussian with 0 mean and a finite parameter $v_{2}^{2}S_{2}>0$.
Assumption 5.6 (Design Matrix).
Consider the random design (whitened) matrices denoted by $\mathbf{Q}=(\mathbf{W}_{t}-\bm{\mu}_{t})\Sigma_{t}^{-1/2}$ and $\mathbf{P}=(\mathbf{X}_{c}-\bm{\mu}_{c})\Sigma_{c}^{-1/2}$ , where $\Sigma_{t}=\operatorname{Var}[\mathbf{W}_{t,i}\mid A_{i}=1]$ is the conditional covariance matrix of $\mathbf{W}_{t}=(\mathbf{X}_{t}\quad\mathbf{M}_{t})$ for the treatment observations, $\Sigma_{c}=\operatorname{Var}\left[X_{i}\mid A_{i}=0\right]$ is the conditional covariance matrix for the control observations, and $\mu_{t}=\mathbb{E}[\mathbf{W}_{t,i}\mid A_{i}=1]$ and $\mu_{c}=\mathbb{E}[X_{i}\mid A_{i}=0]$. We denote $\mu=\mathbb{E}[X]$ and $\tilde{a}=\left[\mu\quad\mathbf{B}_{0}\mu\right]^{\top}$. We suppose the following,
  • $\mathbb{E}[Q_{ij}]=0$, $\mathbb{E}[P_{ij}]=0$ and the variance is unity, $\operatorname{Var}[Q_{ij}]=1$, $\operatorname{Var}[P_{ij}]=1$, applicable for every pair of indices $i,j$.
  • Rows of $\mathbf{Q}$ and $\mathbf{P}$, denoted as $Q_{i}$ and $P_{i}$ respectively are all i.i.d sub-Gaussian random vectors with some sub-Gaussian constants $\iota>0$ and $\varsigma>0$.
  • The diagonal of $\Sigma_{t}$ and $\Sigma_{c}$ are entry-wise bounded, hence $(\Sigma_{t})_{jj}\leqslant S_{1}$ for all $j=1,\ldots,p+q$ and that $(\Sigma_{c})_{jj}\leqslant S_{2}$ for all $j=1,\ldots,p$.
  • The eigenvalue of $\Sigma_{t}$ and $\Sigma_{c}$ are lower and upper bounded by some universal constants $(C_{t,\text{min}},C_{t,\text{max}})$ and $(C_{c,\text{min}},C_{c,\text{max}})$, say $\sigma_{\text{min}}(\Sigma_{t})\geq C_{t,\text{min}}>0$, $\sigma_{\text{min}}(\Sigma_{c})\geq C_{c,\text{min}}>0$, and $\sigma_{\text{max}}(\Sigma_{t})\leqslant C_{t,\text{max}}<\infty$, $\sigma_{\text{max}}(\Sigma_{c})\leqslant C_{c,\text{max}}\leq\infty$.
  • The operator norm $\|\mathbf{B}_{0}\|_{\text{op}}\leqslant C$, $\|\mathbf{B}_{0}\|_{1,\infty}\leq C_{B}$, $\|\mu\|_{\infty}\leqslant C_{X}$ and $\|\gamma_{1}\|_{2}\leqslant C_{\gamma}$ are finite and upper bounded by some constants $C$, $C_{B}$, $C_{X}$ and $C_{\gamma}$.
  • The $\ell_{\infty}$-norm of $(\mu_{t},\mu_{c})$ is bounded, i.e., $\|\mu_{t}\|_{\infty}\leq C_{t}$, $\|\mu_{c}\|_{\infty}\leq C_{c}$ for some constants $C_{t}>0$ and $C_{c}>0$.
  • There exists some constant $\tilde{c}$ and $\kappa$, independent of the sample size $n$ such that $q/p^{\tilde{c}}\rightarrow 0$, and $\tilde{a}^{\top}\Sigma^{-1}_{W}\tilde{a}\leq V_{1}$, $\mu^{\top}\Sigma^{-1}_{c}\mu\leq V_{2}$, i.e., $\tilde{a}^{\top}\Sigma_{t}^{-1}\tilde{a}=\mathcal{O}(1)$, $\mu^{\top}\Sigma_{c}^{-1}\mu=\mathcal{O}(1)$, and the maximum entries of $\tilde{a},\mu$ satisfies $\lVert\tilde{a}\rVert_{\infty},\lVert\mu\rVert_{\infty}\geq\kappa>0$.
Assumption 5.5 and 5.6 are commonly used for analyzing high-dimensional statistical models (e.g., see Bühlmann and Van De Geer, (2011); Lin et al., (2023)). These assumptions impose a restriction on the tail behavior of the errors and the design matrices and facilitate the use of many existing concentration inequalities. In Assumption 5.6, we also assumed that the diagonals of the variance matrices and the operator norm of $\mathbf{B}_{0}$ are finite. The finiteness of the operator norm of $\mathbf{B}_{0}$ implies that the rows of $\mathbf{W}$ are sub-Gaussian, as elaborated in the proof of Theorem 5.11. While it is possible to relax the boundedness assumptions by allowing the bounds to grow with $n$, it would require careful tracking of the constants, as they would affect the estimator’s convergence rate. Since this would not contribute significantly to the core idea of the paper, we choose not to pursue this generalization.
Remark 5.7.
In the setting in our work, it is not reasonable to assume that the coordinates of $\mathbf{W}_{t}$ are independent, as the coordinates of $M_{i}$ depend on each other through $X_{i}$ in our data-generating process. This significantly differs from the theoretical analysis in (Athey et al., , 2018), where the authors assume that individual entries are all independent.
Remark 5.8.
In the proof of the Theorem 5.11, we require a minimum estimand size $\lVert a\rVert_{\infty},\lVert\bar{X}\rVert_{\infty}\geq 2\kappa$. This is required to rule out the fact that if the contrasts are relatively small, then we would end up with having $\tau_{1},\tau_{2}=0$ resulting in a super-efficient estimator as noted in Athey et al., (2018).
Assumption 5.9.
We assume that the population covariance matrices $\Sigma_{t}$ and $\Sigma_{c}$ satisfy the $\{k,\eta_{t},L_{t}\}$- and $\{s,\eta_{c},L_{c}\}$-restricted eigenvalue condition respectively, with its corresponding compatibility constant
$\varphi^{2}_{W}(\Sigma_{t},\eta_{t},L_{t})$ and $\varphi_{c}^{2}(\Sigma_{c},\eta_{c},L_{c})$ respectively (see Definition 5.3 for details).
The Restricted Eigenvalue (RE) condition addresses the limitations of relying solely on the minimum eigenvalue being bounded away from zero by ensuring that the design matrix behaves well on sparse subsets of predictors. This condition is essential for consistently estimating the (sparse) parameter of interest in high-dimensional settings, as elaborated in Bickel et al., (2009) and Hastie et al., (2015). Moreover, Zhang et al., (2014) indicates that without the RE condition, no polynomial-time computable estimator can achieve the optimal convergence rates for sparse linear regression.
5.2 Consistency and Asymptotic Normality of $\hat{\theta}_{0}$
In this subsection, we present our key theoretical finding, i.e., our estimator $\hat{\theta}_{0}$ is consistent and asymptotically normal. As stated in Algorithm 1, $\hat{\theta}_{0}$ can be written as
$\hat{\theta}_{0}$ $=\bar{X}^{\top}\left(\hat{\beta}_{1}+\hat{b}\right)+\sum_{\{\begin{subarray}{c}i:A_{i}=1\end{subarray}\}}\tau_{1,i}\left(Y_{i}-\mathbf{W}_{t,i}\cdot\hat{\phi}_{1}\right)+\sum_{\{i:A_{i}=0\}}\tau_{2,i}\left(M_{i}^{\top}\hat{\gamma}_{1}-X_{i}^{\top}\hat{b}\right)\,,$ (5.3)
where
$\tau_{1}$ $=\operatorname{argmin}_{\tilde{\tau}_{1}}\left\{\left\lVert\tilde{\tau}_{1}\right\rVert_{2}^{2}\quad\text{subject to}\quad\left\lVert a-\mathbf{W}_{t}^{\top}\tilde{\tau}_{1}\right\rVert_{\infty}\leqslant K_{1}\sqrt{\frac{\log(p+q)}{n_{t}}},\|\tilde{\tau}_{1}\|_{\infty}\leq n_{t}^{-2/3}\right\}\,,$ (5.4)
$\tau_{2}$ $=\operatorname{argmin}_{\tilde{\tau}_{2}}\left\{\|\tilde{\tau}_{2}\|_{2}^{2}\quad\text{ subject to }\quad\|\bar{X}-\mathbf{X}_{c}^{\top}\tilde{\tau}_{2}\|_{\infty}\leqslant K_{2}\sqrt{\frac{\log(p)}{n_{c}}},\|\tilde{\tau}_{2}\|_{\infty}\leq n_{c}^{-2/3}\right\}\,,$ (5.5)
with $a=\begin{bmatrix}\bar{X}&(\hat{\mathbf{B}}_{0}\bar{X})\end{bmatrix}^{\top}$, $\hat{\phi}_{1}$ and $\hat{b}$ are sparse linear estimator of $\phi_{1}$ and $b$, respectively, and $K_{1}$ and $K_{2}$ are hyper-parameters (uniformly bounded in $(n,p)$). Equations (5.4) and (5.5) are optimization problems designed to obtain the weights for the bias correction terms in the estimator $\hat{\theta}_{0}$. Thus, it is crucial to establish that these optimization problems are feasible (with high probability), a result we formalize in the following lemma.
Lemma 5.10.
Under Assumption 5.6, the optimization problems (5.4) and (5.5) are feasible with probability going 1 for any choice of $K_{1}\geq C\iota^{2}\sqrt{V_{1}^{\prime}S_{1}}$, and $K_{2}\geq C\mathcal{\varsigma}^{2}\sqrt{V_{2}S_{2}}$.
The next result outlines the asymptotic normality of our estimator $\hat{\theta}_{0}$, based on previous assumptions and the Lemma 5.10.
Theorem 5.11.
Under Assumptions 5.4, 5.5, 5.6, 5.9, the specification of $(K_{1},K_{2})$ in Lemma 5.10, and the choice $\lambda_{1,n}=\iota^{2}v_{1}\sqrt{(c_{1}^{2}\log(p+q))/n_{t}}$, $\lambda_{2,n}=\varsigma^{2}v_{2}\sqrt{(c_{2}^{2}\log(p))/n_{c}}$, our estimator $\hat{\theta}_{0}$ satisfies:
$\frac{\sqrt{n}(\hat{\theta}_{0}-\theta_{0})}{\sigma_{n}}\xrightarrow{\mathcal{D}}\mathcal{N}(0,1)$ (5.6)
where $\sigma_{n}^{2}=n\left(\sigma_{1}^{2}\|\tau_{1}\|_{2}^{2}+\sigma_{2}^{2}\|\tau_{2}\|_{2}^{2}\|\gamma_{1}\|_{2}^{2}\right)$ with $\sigma_{1}^{2},\sigma_{2}^{2}$ are as defined in Assumption 5.5.
Remark 5.12.
When $\gamma_{1}=0$, i.e., the response is unaffected by the mediators, then the Lasso estimator $\hat{\gamma}_{1}$ will be equal to 0 with high probability. In that case, Theorem 5.11 implies $(\hat{\theta}_{0}-\theta_{0})/(\sigma_{1}\|\tau_{1}\|_{2})$ is asymptotically normal, which is Theorem 3 of Athey et al., (2018). Therefore, our result can be viewed as a generalized version of the results in (Athey et al., , 2018) in the presence of high dimensional mediators along with the high dimensional covariates.
6 Conclusion
In this paper, we introduced a novel methodology for estimating the mediation functional in the presence of high-dimensional covariates and mediators, potentially exceeding the sample size. Our approach allows for the covariate-treatment interaction in the generation of the mediators, as well as both covariate-treatment and mediator-treatment interactions in the generation of the response. The main contribution of our work is a novel debiasing technique in the presence of these two challenges. The structure of the mediation functional, which was our parameter of interest, possesses certain properties which required development of new debiasing methodology. Our proposed methodology can be of independent interest in other statistical settings where the parameter of interest has a structure similar to that of the mediation functional. We showed that our estimator is root-$n$ consistent and asymptotically normal, which aids in performing statistical inference on the direct/indirect treatment effects by providing an asymptotically valid confidence interval. To the best of our knowledge, this is the first study to address the challenge of estimating NDE and NIE when dealing with high-dimensional covariates, mediators, and their interactions with the treatment variable simultaneously. While this paper focuses on a linear model, extending the methodology to generalized linear models is a natural next step, which we leave as an intriguing direction for future research.
Appendix A Proofs
To simplify the notation, we assume there are $2n_{t}$ observations in the treatment group and $2n_{c}$ observations in the control group. Since we are performing data splitting (Algorithm 1), each half of the data consists of $n_{t}$ treatment samples and $n_{c}$ control samples.
A.1 Proof of Proposition 4.1
We begin our analysis by using the contrast $a=\begin{bmatrix}\bar{X}&(\mathbf{\hat{B}}_{0}\bar{X})\end{bmatrix}$ and $b=\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}$, define
$\hat{\vartheta}$ $=a^{\top}\hat{\phi}_{1}+\tau_{1}^{\top}(Y_{t}-\mathbf{X}_{t}^{\top}\hat{\beta}_{1}-\mathbf{M}_{t}^{\top}\hat{\gamma}_{1})+\bar{X}^{\top}\hat{b}+\tau_{2}^{\top}(\mathbf{M}_{c}^{\top}\hat{\gamma}_{1}-\mathbf{X}_{c}^{\top}\hat{b})$
$=a^{\top}\hat{\phi}_{1}+\tau_{1}^{\top}(Y_{t}-\mathbf{W}_{t}\cdot\hat{\phi}_{1})+\bar{X}^{\top}\hat{b}+\tau_{2}^{\top}(\mathbf{M}_{c}^{\top}\hat{\gamma}_{1}-\mathbf{X}_{c}^{\top}\hat{b})\,.$
Dealing with the first part, we have:
$a^{\top}\hat{\phi}_{1}+\tau_{1}^{\top}(Y_{t}-\mathbf{W}_{t}\cdot\hat{\phi}_{1})$
$=a^{\top}\phi_{1}+a^{\top}\left(\hat{\phi}_{1}-\phi_{1}\right)+\tau_{1}^{\top}(Y_{t}-\mathbf{W}_{t}\cdot{\phi}_{1})+\tau_{1}^{\top}\mathbf{W}_{t}\left(\phi_{1}-\hat{\phi}_{1}\right)$
$=a^{\top}\phi_{1}+\left(a-\mathbf{W}_{t}^{\top}\tau_{1}\right)^{\top}\left(\hat{\phi}_{1}-\phi_{1}\right)+\tau_{1}^{\top}\epsilon$
$=\bar{X}^{\top}\beta_{1}+\left(\hat{\mathbf{B}}_{0}\bar{X}\right)^{\top}\gamma_{1}+\left(a-\mathbf{W}_{t}^{\top}\tau_{1}\right)^{\top}\left(\hat{\phi}_{1}-\phi_{1}\right)+\tau_{1}^{\top}\epsilon\,.$
Then the second part, we obtain:
$\bar{X}^{\top}\hat{b}+\tau_{2}^{\top}\left(\mathbf{M}_{c}^{\top}\hat{\gamma}_{1}-\mathbf{X}_{c}^{\top}\hat{b}\right)$
$=\bar{X}^{\top}b+\bar{X}^{\top}(\hat{b}-b)+\tau_{2}^{\top}\left(\mathbf{M}_{c}^{\top}\hat{\gamma}_{1}-\mathbf{X}_{c}^{\top}b+\mathbf{X}_{c}^{\top}(b-\hat{b})\right)$
$=\bar{X}^{\top}b+\left(\bar{X}-\mathbf{X}_{c}^{\top}\tau_{2}\right)^{\top}(\hat{b}-b)+\tau_{2}^{\top}U^{\top}\hat{\gamma}_{1}$
$=\bar{X}^{\top}\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}+\left(\bar{X}-\mathbf{X}_{c}^{\top}\tau_{2}\right)^{\top}(\hat{b}-b)+\tau_{2}^{\top}U^{\top}\hat{\gamma}_{1}\,.$
By linearity,
$\hat{\vartheta}=\bar{X}^{\top}\beta_{1}+\left(\hat{\mathbf{B}}_{0}\bar{X}\right)^{\top}\gamma_{1}+\bar{X}^{\top}\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}+\left(a-\mathbf{W}_{t}^{\top}\tau_{1}\right)^{\top}\left(\hat{\phi}_{1}-\phi_{1}\right)+\left(\bar{X}-\mathbf{X}_{c}^{\top}\tau_{2}\right)^{\top}(\hat{b}-b)+\tau_{1}^{\top}\epsilon+\tau_{2}^{\top}U^{\top}\hat{\gamma}_{1}\,.$
Hence, we can achieve the product forms by subtracting $(\hat{\mathbf{B}}_{0}\bar{X})^{\top}\hat{\gamma}_{1}$, and our estimator becomes
$\hat{\theta}_{0}=\hat{\vartheta}-(\hat{\mathbf{B}}_{0}\bar{X})^{\top}\hat{\gamma}_{1}$
$=\bar{X}^{\top}\beta_{1}+\left(\hat{\mathbf{B}}_{0}\bar{X}\right)^{\top}\left(\gamma_{1}-\hat{\gamma}_{1}\right)+\left(\mathbf{B}_{0}\bar{X}\right)^{\top}\hat{\gamma}_{1}+\left(a-\mathbf{W}_{t}^{\top}\tau_{1}\right)^{\top}\left(\hat{\phi}_{1}-\phi_{1}\right)+\left(\bar{X}-\mathbf{X}_{c}^{\top}\tau_{2}\right)^{\top}(\hat{b}-b)+\tau_{1}^{\top}\epsilon+\tau_{2}^{\top}U^{\top}\hat{\gamma}_{1}$
$=\bar{X}^{\top}\beta_{1}+\left(\mathbf{B}_{0}\bar{X}\right)^{\top}\gamma_{1}+\left(\hat{\mathbf{B}}_{0}\bar{X}\right)^{\top}\left(\gamma_{1}-\hat{\gamma}_{1}\right)+\left(\mathbf{B}_{0}\bar{X}\right)^{\top}\left(\hat{\gamma}_{1}-\gamma_{1}\right)+\left(a-\mathbf{W}_{t}^{\top}\tau_{1}\right)^{\top}\left(\hat{\phi}_{1}-\phi_{1}\right)+\left(\bar{X}-\mathbf{X}_{c}^{\top}\tau_{2}\right)^{\top}\left(\hat{b}-b\right)$
$~{}~{}~{}~{}~{}~{}~{}+\tau_{1}^{\top}\epsilon+\tau_{2}^{\top}U^{\top}\hat{\gamma}_{1}$
$=\theta_{0}+\left(\hat{\mathbf{B}}_{0}\bar{X}-\mathbf{B}_{0}\bar{X}\right)^{\top}\left(\gamma_{1}-\hat{\gamma}_{1}\right)+\left(a-\mathbf{W}_{t}^{\top}\tau_{1}\right)^{\top}\left(\hat{\phi}_{1}-\phi_{1}\right)+\left(\bar{X}-\mathbf{X}_{c}^{\top}\tau_{2}\right)^{\top}\left(\hat{b}-b\right)+\tau_{1}^{\top}\epsilon+\tau_{2}^{\top}U^{\top}\hat{\gamma}_{1}\,.$
We move the $\theta_{0}$ to the left hand side to get
$\hat{\theta}_{0}-\theta_{0}=\left(\hat{\mathbf{B}}_{0}\bar{X}-\mathbf{B}_{0}\bar{X}\right)^{\top}\left(\gamma_{1}-\hat{\gamma}_{1}\right)+\left(a-\mathbf{W}_{t}^{\top}\tau_{1}\right)^{\top}\left(\hat{\phi}_{1}-\phi_{1}\right)+\left(\bar{X}-\mathbf{X}_{c}^{\top}\tau_{2}\right)^{\top}\left(\hat{b}-b\right)+\tau_{1}^{\top}\epsilon+\tau_{2}^{\top}U^{\top}\hat{\gamma}_{1}\,,$
where
$V_{n}=\tau_{1}^{\top}\epsilon+\tau_{2}^{\top}U^{\top}\hat{\gamma}_{1}$
$\Delta_{n}=\left(\hat{\mathbf{B}}_{0}\bar{X}-\mathbf{B}_{0}\bar{X}\right)^{\top}\left(\gamma_{1}-\hat{\gamma}_{1}\right)+\left(a-\mathbf{W}_{t}^{\top}\tau_{1}\right)^{\top}\left(\hat{\phi}_{1}-\phi_{1}\right)+\left(\bar{X}-\mathbf{X}_{c}^{\top}\tau_{2}\right)^{\top}\left(\hat{b}-b\right)\,,$
the proof is concluded by the Hölder’s inequality. $\square$
A.2 Proof of Lemma 5.10
We prove Lemma 5.10 by constructing $\tau_{1}^{*}$ and $\tau_{2}^{*}$, which are feasible solutions of optimization problems (5.4) and (5.5) respectively with high probability, where:
$\tau_{1,i}^{*}=\frac{1}{n_{t}}\check{a}^{\top}\Sigma_{t}^{-1}(\mathbf{W}_{t,i}-\mu_{t}),\qquad\tau_{2,i}^{*}=\frac{1}{n_{c}}\mu^{\top}\Sigma_{c}^{-1}(\mathbf{X}_{c,i}-\mu_{c}),\qquad 1\leq i\leq n\,.$ (A.1)
where $\check{a}=[\mu\quad\hat{\mathbf{B}}_{0}\mu],\mu=\mathbb{E}[X]$. We start with $\tau_{1}^{*}$ and show that with high probability, it satisfies the constraints of Equation (5.4), i.e.
$\|a-\mathbf{W}_{t}^{\top}\tau_{1}^{*}\|_{\infty}\leq K_{1}\sqrt{\log(p+q)/n_{t}}\,,\quad\|\tau_{1}^{*}\|_{\infty}\leq n_{t}^{-2/3}\,.$ (A.2)
for some constant $K_{1}>0$ to be specified in the proof. Towards that direction, first, observe that:
$\|a-\mathbf{W}_{t}^{\top}\tau_{1}^{*}\|_{\infty}$ $\leq\|a-\check{a}\|_{\infty}+\|\check{a}-\mathbf{W}_{t}^{\top}\tau_{1}^{*}\|_{\infty}\,.$ (A.3)
For the first term on the right-hand side of (A.3),
$\|a-\check{a}\|_{\infty}=\|[\bar{X}\quad\hat{\mathbf{B}}_{0}\bar{X}]-[\mu\quad\hat{\mathbf{B}}_{0}\mu]\|_{\infty}=\max\{\|\bar{X}-\mu\|_{\infty},\|\hat{\mathbf{B}}_{0}(\bar{X}-\mu)\|_{\infty}\}\,.$ (A.4)
As we have assumed $X_{i}$’s sub-Gaussian (Assumption 5.6), a simple application of Hoeffding’s inequality along with a union bound yields $\|\bar{X}-\mu\|_{\infty}\leq K_{2}\sqrt{\log{(p)}/n}$ with probability approaching to 1. For the other term, we use the fact that with probability going to 1, $\|\hat{\mathbf{B}}_{0}-\mathbf{B}_{0}\|_{1,\infty}\leq C_{1}s\sqrt{\log(p)/n_{c}}$ for some constant $C_{1}>0$, which is later established in Lemma A.1. Using this information, we bound the $\|\hat{\mathbf{B}}_{0}(\bar{X}-\mu)\|_{\infty}$ as follows:
$\|\hat{\mathbf{B}}_{0}\bar{X}-\hat{\mathbf{B}}_{0}\mu\|_{\infty}$ $=\|\mathbf{B}_{0}\bar{X}-\mathbf{B}_{0}\mu-\hat{\mathbf{B}}_{0}\bar{X}-\hat{\mathbf{B}}_{0}\mu-\mathbf{B}_{0}\bar{X}+\mathbf{B}_{0}\mu\|_{\infty}$
$\leq\|\mathbf{B}_{0}(\bar{X}-\mu)\|_{\infty}+\|\hat{\mathbf{B}}_{0}-\mathbf{B}_{0}\|_{1,\infty}\|\bar{X}-\mu\|_{\infty}$
$=\max_{1\leq j\leq q}\left|\left\langle\mathbf{B}_{0,j\cdot},\bar{X}-\mu\right\rangle\right|+K_{2}C_{1}s\frac{\log{(p)}}{n_{c}}$
$\leq\max_{1\leq j\leq q}\|\mathbf{B}_{0,j\cdot}\|_{1}\|\bar{X}-\mu\|_{\infty}+K_{2}C_{1}s\frac{\log{(p)}}{n_{c}}$
$\leq 2\|\mathbf{B}_{0}\|_{1,\infty}\|\bar{X}-\mu\|_{\infty}+K_{2}C_{1}s\frac{\log{(p)}}{n_{c}}$
$\leq C_{B}K_{2}\sqrt{\frac{\log{(p})}{n}}+K_{2}C_{1}s\frac{\log{(p)}}{n_{c}}:=K_{3}\sqrt{\frac{\log{(p})}{n_{c}}}\,,$
where the last inequality follows $\|\mathbf{B}_{0}\|_{1,\infty}\leq C_{B}$ (Assumption 5.6) and $s\sqrt{\log{(p)}/n_{c}}=o(1)$ (Assumption 5.4). Thus, we conclude that with probability going to 1:
$\|a-\check{a}\|_{\infty}\leq(K_{2}\vee K_{3})\sqrt{\frac{\log{(p})}{n_{c}}}\,.$ (A.5)
Next, we provide an upper bound on $\|\check{a}-\mathbf{W}_{t}^{\top}\tau_{1}^{*}\|_{\infty}$ in Equation (A.4). Towards that end, we write $(\mathbf{W}_{t}^{\top}\tau_{1}^{*})$ as:
$\mathbf{W}_{t}^{\top}\tau_{1}^{*}$ =$=\sum_{\{i:A_{i}=1\}}\mathbf{W}_{t,i}\tau_{1,i}^{*}=\frac{1}{n_{t}}\sum_{i}\mathbf{W}_{t,i}(\mathbf{W}_{t,i}-\mu_{t})^{\top}\Sigma_{t}^{-1}\check{a}$
$=\underbrace{\frac{1}{n_{t}}\sum_{\{i:A_{i}=1\}}(\mathbf{W}_{t,i}-\mu_{t})(\mathbf{W}_{t,i}-\mu_{t})^{\top}\Sigma_{t}^{-1}\check{a}}_{:=\xi_{1}}+\underbrace{\mu_{t}\frac{1}{n_{t}}\sum_{i}(\mathbf{W}_{t,i}-\mu_{t})^{\top}\Sigma_{t}^{-1}\check{a}}_{:=\xi_{2}}\,.$ (A.6)
For any $1\leq j\leq p+q$, we can simplify the coordinates $\xi_{1,j}$ as follows:
$\xi_{1,j}$ $=\frac{1}{n_{t}}e_{j}^{\top}\sum_{\{i:A_{i}=1\}}(\mathbf{W}_{t,i}-\mu_{t})(\mathbf{W}_{t,i}-\mu_{t})^{\top}\Sigma_{t}^{-1}\check{a}$
$=\frac{1}{n_{t}}e_{j}^{\top}\sum_{\{i:A_{i}=1\}}(\Sigma_{t}^{1/2}Q_{i})(\Sigma_{t}^{1/2}Q_{i})^{\top}\Sigma_{t}^{-1}\check{a}$
$=\frac{1}{n_{t}}\sum_{\{i:A_{i}=1\}}(e_{j}^{\top}\Sigma_{t}^{1/2}Q_{i})(Q_{i}^{\top}\Sigma_{t}^{-1/2}\check{a})$
$=\frac{1}{n_{t}}\sum_{\{i:A_{i}=1\}}(Q_{i}^{\top}\Sigma_{t}^{-1/2}\check{a})(e_{j}^{\top}\Sigma_{t}^{1/2}Q_{i})$
$=\frac{1}{n_{t}}\sum_{\{i:A_{i}=1\}}Q_{i}^{\top}A_{j}Q_{i},\quad A_{j}:=\Sigma_{t}^{-\frac{1}{2}}\check{a}e_{j}^{\top}\Sigma_{t}^{\frac{1}{2}}\,,$
where $Q_{i}$ denote the $i$-th row of the whitened matrix $\mathbf{Q}$ (as defined in Assumption 5.6), and $e_{j}$ is the $j$-th canonical basis vector. Using this, we have:
$\|\check{a}-\mathbf{W}_{t}^{\top}\tau_{1}^{*}\|_{\infty}\leq\|\check{a}-\xi_{1}\|_{\infty}+\|\xi_{2}\|_{\infty}=\max_{1\leq j\leq p+q}\left|\check{a}_{j}-\frac{1}{n_{t}}\sum_{i:A_{i}=1}Q_{i}^{\top}A_{j}Q_{i}\right|+\|\xi_{2}\|_{\infty}\,.$ (A.7)
We will next focus on upper bounding the first part of the right-hand side of Equation (A.7). Observe that, as $Q_{i}$’s are sub-Gaussian (Assumption 5.5), the random variables $Q_{i}^{\top}A_{j}Q_{i}$ are sub-Exponential. For all $1\leq j\leq p+q$, the matrix $A_{j}$ is a rank 1 matrix and its Frobenius norm is uniformly upper bounded as follows:
$\left\|A_{j}\right\|_{F}^{2}=\operatorname{tr}\left(A_{j}^{\top}A_{j}\right)$ $=\operatorname{tr}\left(\Sigma_{t}^{\frac{1}{2}}e_{j}\check{a}^{\top}\Sigma_{W}^{-\frac{1}{2}}\Sigma_{W}^{-\frac{1}{2}}\check{a}e_{j}^{\top}\Sigma_{W}^{\frac{1}{2}}\right)$
$=\operatorname{tr}\left(\Sigma_{W}^{\frac{1}{2}}e_{j}\check{a}^{\top}\Sigma_{W}^{-1}\check{a}e_{j}^{\top}\Sigma_{W}^{\frac{1}{2}}\right)$
$=\operatorname{tr}\left(e_{j}\Sigma_{W}^{\frac{1}{2}}\Sigma_{t}^{\frac{1}{2}}\check{a}^{\top}\Sigma_{W}^{-1}\check{a}\right)$
$=\left(\Sigma_{W}\right)_{jj}\check{a}^{\top}\Sigma^{-1}_{W}\check{a}\leqslant V_{1}^{\prime}S_{1}\,,$
where the bound in last line follows form Assumption 5.6 and Lemma A.1. Furthermore, observe that conditionally on treatment indicators and the observations in $\mathcal{D}_{1}$ (call this sigma field $\mathcal{F}_{0,n}$), $Q_{i}^{\top}A_{j}Q_{i}$ (for $i$ with $A_{i}=1$) are i.i.d with mean $\check{a}_{j}$. Hence, an application of Corollary 2.8 of Zajkowski, (2020) (see Lemma A.2 for details) yields:
$\mathbb{E}\left[{\sf exp}\left\{t\left(Q_{i}^{\top}A_{j}Q_{i}-\check{a}_{j}\right)\right\}\mid\mathcal{F}_{0,n}\right]\leqslant{\sf exp}\left\{8C_{1}^{2}t^{2}\iota^{4}V_{1}^{\prime}S_{1}\right\},\quad\forall t\leqslant\frac{C_{2}}{2\iota^{2}\sqrt{2V_{1}^{\prime}S_{1}}}.$ (A.8)
This implies for all $t\leq\sqrt{n_{t}}C_{2}/(2\iota^{2}\sqrt{2V_{1}^{\prime}S_{1}})$, we have:
$\mathbb{E}\left[{\sf exp}\left\{\sqrt{n_{t}}t\left(\xi_{1}-\check{a}\right)_{j}\right\}\mid\mathcal{F}_{0,n}\right]$ $=\mathbb{E}\left[{\sf exp}\left\{\frac{t}{\sqrt{n_{t}}}\sum_{\{i:A_{i}=1\}}(Q_{i}^{\top}A_{j}Q_{i}-\check{a}_{j})\right\}\mid\mathcal{F}_{0,n}\right]$
=$=\prod_{\{i:A_{i}=1\}}\mathbb{E}\left[{\sf exp}\left\{\frac{t}{\sqrt{n_{t}}}(Q_{i}^{\top}A_{j}Q_{i}-\check{a}_{j})\right\}\mid\mathcal{F}_{0,n}\right]$
$\leq\prod_{\{i:A_{i}=1\}}{\sf exp}\left\{8C_{1}^{2}(t^{2}/n_{t})\iota^{4}V_{1}^{\prime}S_{1}\right\}$
$={\sf exp}\left\{8C_{1}^{2}t^{2}\iota^{4}V_{1}^{\prime}S_{1}\right\}\,.$ (A.9)
Applying Chernoff’s inequality, we conclude:
$\mathbb{P}\left(\sqrt{n_{t}}\|\xi_{1}-\check{a}\|_{\infty}\geq 9C_{1}^{2}t\iota^{4}V_{1}^{\prime}S_{1}\mid\mathcal{F}_{0,n}\right)$ $\leq\sum_{j=1}^{p+q}\mathbb{P}\left(\sqrt{n_{t}}\left(\xi_{1}-\check{a}\right)_{j}\geqslant 9C_{1}^{2}t\iota^{4}V_{1}^{\prime}S_{1}\mid\mathcal{F}_{0,n}\right)$
$=\sum_{j=1}^{p+q}\mathbb{P}\left(\sqrt{n_{t}}t\left(\xi_{1}-\check{a}\right)_{j}\geqslant 9C_{1}^{2}t^{2}\iota^{4}V_{1}^{\prime}S_{1}\mid\mathcal{F}_{0,n}\right)$
$=\sum_{j=1}^{p+q}\mathbb{P}\left({\sf exp}{(\sqrt{n_{t}}t\left(\xi_{1}-\check{a}\right)_{j})}\geq{\sf exp}{(9C_{1}^{2}t^{2}\iota^{4}V_{1}^{\prime}S_{1})}\mid\mathcal{F}_{0,n}\right)$
$\leq\sum_{j=1}^{p+q}{\sf exp}{(-9C_{1}^{2}t^{2}\iota^{4}V_{1}^{\prime}S_{1})}\mathbb{E}\left[{\sf exp}{(\sqrt{n_{t}}t\left(\xi_{1}-\check{a}\right)_{j})}\right]$
$={\sf exp}{(\log{(p+q)}-C_{1}^{2}t^{2}\iota^{4}V_{1}^{\prime}S_{1})}\,.$ (A.10)
The above inequality is value for any choice of $t\leq\sqrt{n_{t}}C_{2}/(2\iota^{2}\sqrt{2V_{1}^{\prime}S_{1}})$. If we choose $t=\sqrt{2\log{(p+q)}}/\iota^{2}\sqrt{C_{1}^{2}V_{1}^{\prime}S_{1}}$ (which is valid as $\log{(p+q)}\ll\sqrt{n}$), we have from Equation (A.2):
$\mathbb{P}\left(\sqrt{n_{t}}\|\xi_{1}-\check{a}\|_{\infty}\geq 9C_{1}^{2}t\iota^{4}V_{1}^{\prime}S_{1}\mid\mathcal{F}_{0,n}\right)$ $=\mathbb{P}\left(\|\xi_{1}-\check{a}\|_{\infty}\geq 9C_{1}\iota^{2}\sqrt{2V_{1}^{\prime}S_{1}}\sqrt{\frac{\log{(p+q)}}{n_{t}}}\mid\mathcal{F}_{0,n}\right)$
$\leq(p+q)^{-1}\,.$ (A.11)
Defining $9C_{1}\iota^{2}\sqrt{2V_{1}S_{1}}=K_{4}$ and taking expectation with respect to the $\sigma$-field $\mathcal{F}_{0,n}$ on the both side of Equation (A.2), we conclude that $\|\xi_{1}-\check{a}\|_{\infty}\leq K_{4}\sqrt{\log{(p+q)}/n_{t}}$ with probability approaching to 1.
Next, we provide an upper bound on $\|\xi_{2}\|_{\infty}$. From the definition of $\xi_{2}$ (Equation (A.2)) we have:
$\|\xi_{2}\|_{\infty}\leq\left\|\mu_{t}\right\|_{\infty}\left|\frac{1}{n_{t}}\sum_{\{i:A_{i}=1\}}(\mathbf{W}_{t,i}-\mu_{t})^{\top}\Sigma_{t}^{-1}\check{a}\right|=\left\|\mu_{t}\right\|_{\infty}\left|\frac{1}{n_{t}}\sum_{\{i:A_{i}=1\}}Q_{i}^{\top}\Sigma_{t}^{-1/2}\check{a}\right|:=\left\|\mu_{t}\right\|_{\infty}\left|\frac{1}{n_{t}}\sum_{\{i:A_{i}=1\}}Z_{t,i}\right|\,,$
where $Z_{t,i}=Q_{i}^{\top}\Sigma_{t}^{-1/2}\check{a}$. Note that $Z_{t,i}\in\mathbb{R}$ and therefore, by Chebychev’s inequality, we have
$\mathbb{P}\left(\left|\frac{1}{n_{t}}\sum_{\{i:A_{i}=1\}}Z_{t,i}\right|\geq t\mid\mathcal{F}_{0,n}\right)\leq\frac{\operatorname{Var}\left(Z_{t}\mid\mathcal{F}_{0,n}\right)}{n_{t}t^{2}}\,.$
Now $\operatorname{Var}(Z_{t}\mid A=1)=\check{a}^{\top}\Sigma_{t}^{-1}\check{a}$ which is upper bounded by $V_{1}^{\prime}$ by Lemma A.1. Therefore, taking $t=\sqrt{\log{(p+q)}/n_{t}}$ we obtain
$\mathbb{P}\left(\left|\frac{1}{n_{t}}\sum_{i:A_{i}=1}Z_{t,i}\right|\geq\sqrt{\frac{\log{(p+q)}}{n_{t}}}\mid\mathcal{F}_{0,n}\right)\leq\frac{\operatorname{Var}\left(Z_{t}\mid\mathcal{F}_{0,n}\right)}{\log{(p+q)}}\,.$
Furthermore, $\|\mu_{t}\|_{\infty}\leq C_{t}$ by Assumption 5.6. Hence, we conclude that $\|\xi_{2}\|_{\infty}\leq K_{5}\sqrt{\log{(p+q)}/n_{t}}$ with probability going to 1. Finally, combining this bound along with the upper bounds in Equation (A.3) and (A.7) we conclude that $\|a-\mathbf{W}_{t}^{\top}\tau_{1}^{*}\|_{\infty}\leq K_{1}\sqrt{\log{(p+q)}/n_{t}}$ with probability going to 1, for some constant $K_{1}>0$.
Next, we show the second part of the feasibility of the optimization problem (5.4), that is $\|\tau_{1}^{*}\|_{\infty}\leq n_{t}^{-2/3}$. Let $z=\Sigma_{t}^{-1/2}\check{a}$, then $\|z\|_{2}\leq V_{1}^{{}^{\prime}1/2}$ by Lemma A.1. Let $\mathbf{z}=z/\|z\|_{2}$, and as $Q_{i}$’s are sub-Gaussian and by our design (Assumption 5.6), we can write:
[…]
Here $V_{1}^{\prime}$ is a constant as presented in Lemma A.1, and thus by sub-Gaussianity of $Q_{i}$, we have:
[…]
$\leq Ce^{\log n_{t}-\frac{V_{1}^{\prime}n_{t}^{2/3}}{2\iota^{2}}}\,.$
Hence, the optimization problem (5.4) is feasible with high probability with $\|\tau_{1}^{*}\|_{\infty}\leqslant n_{t}^{-2/3}$ and $\|a-\mathbf{W}_{t}^{\top}\tau_{1}^{*}\|_{\infty}\leqslant K_{1}\sqrt{\log(p+q)/n_{t}}$.
We now demonstrate that $\tau_{2,i}^{*}=\frac{1}{n_{c}}\mu^{\top}\Sigma_{c}^{-1}(\mathbf{X}_{c,i}-\mu_{c})$ for $1\leq i\leq n$ is a feasible solution to the optimization problem (5.5) with high probability, i.e.,
$\|\bar{X}-\mathbf{X}_{c}^{\top}\tau_{2}^{*}\|_{\infty}\leqslant K_{5}\sqrt{\log(p)/n_{c}}\,,\ \ \|\tau_{2}^{*}\|_{\infty}\leq n_{c}^{-2/3}\,.$ (A.12)
for some constant $K_{5}>0$ to be specified in the proof. We start with the following decomposition:
$\|\bar{X}-\mathbf{X}_{c}^{\top}\tau_{2}^{*}\|_{\infty}\leq\|\bar{X}-\mu\|_{\infty}+\|\mu-\mathbf{X}_{c}^{\top}\tau_{2}^{*}\|_{\infty}\,.$ (A.13)
For the first term, we have $\|\bar{X}-\mu\|_{\infty}=K_{6}\sqrt{\log{p}/n_{c}}$ with probability going to 1 for some constant $K_{6}>0$, by a simple application of Hoeffding’s inequality along with a union bound. We next obtain the upper bound of $\|\mu-\mathbf{X}_{c}^{\top}\tau_{2}^{*}\|_{\infty}$ in Equation (A.13). Write $(\mathbf{X}_{c}^{\top}\tau_{2}^{*})$ as:
$\mathbf{X}_{c}^{\top}\tau_{2}^{*}$ =$=\sum_{\{i:A_{i}=0\}}X_{c,i}\tau_{2,i}^{*}=\frac{1}{n_{c}}\sum_{\{i:A_{i}=0\}}X_{c,i}^{\top}(X_{c,i}-\mu_{c})\Sigma_{c}^{-1}\mu$
$=\underbrace{\frac{1}{n_{c}}\sum_{\{i:A_{i}=0\}}(X_{c,i}-\mu_{c})(X_{c,i}-\mu_{c})^{\top}\Sigma_{c}^{-1}\mu}_{:=\xi_{3}}+\underbrace{\mu_{c}\frac{1}{n_{c}}\sum_{\{i:A_{i}=0\}}(X_{c,i}-\mu_{c})^{\top}\Sigma_{c}^{-1}\mu}_{:=\xi_{4}}\,.$ (A.14)
Note that for any $1\leq j\leq p$, the $j$-th coordinate of $\xi_{3}$ follows:
$\xi_{3,j}$ $=\frac{1}{n_{c}}e_{j}^{\top}\sum_{\{i:A_{i}=0\}}(X_{c,i}-\mu_{c})(X_{c,i}-\mu_{c})^{\top}\Sigma_{c}^{-1}\mu$
$=\frac{1}{n_{c}}e_{j}^{\top}\sum_{\{i:A_{i}=0\}}(\Sigma_{c}^{1/2}P_{i})(\Sigma_{c}^{1/2}P_{i})^{\top}\Sigma_{c}^{-1}\mu$
$=\frac{1}{n_{c}}\sum_{\{i:A_{i}=0\}}(e_{j}^{\top}\Sigma_{c}^{1/2}P_{i})(P_{i}^{\top}\Sigma_{c}^{-1/2}\mu)$
$=\frac{1}{n_{c}}\sum_{\{i:A_{i}=0\}}(P_{i}^{\top}\Sigma_{c}^{-1/2}\mu)(e_{j}^{\top}\Sigma_{c}^{1/2}P_{i})$
$=\frac{1}{n_{c}}\sum_{\{i:A_{i}=0\}}P_{i}^{\top}G_{j}P_{i},\quad G_{j}=\Sigma_{c}^{-\frac{1}{2}}\mu e_{j}^{\top}\Sigma_{c}^{\frac{1}{2}}\,.$
where $P_{i}$ denote the $i$-th row of the whitened matrix $\mathbf{P}$ (as defined in Assumption 5.6). Applying this, we obtain:
$\|\mu-\mathbf{X}_{c}^{\top}\tau_{2}^{*}\|_{\infty}\leq\|\mu-\xi_{3}\|_{\infty}+\|\xi_{4}\|_{\infty}=\max_{1\leq j\leq p}\left|\mu_{j}-\frac{1}{n_{c}}\sum_{i:A_{i}=0}P_{i}^{\top}G_{j}P_{i}\right|+\|\xi_{4}\|_{\infty}\,.$ (A.15)
Next, we will concentrate on deriving an upper bound for the first term on the right-hand side of Equation (A.15). Similarly, as $P_{i}$’s are sub-Gaussian (Assumption 5.6), the random variables $P_{i}^{\top}G_{j}P_{i}$ are sub-Exponential. Then, the Frobenius norm of $G_{j}$ is a rank-1 matrix, for all $1\leq j\leq p$, can be upper bounded as follows:
$\|G_{j}\|_{F}^{2}=tr\left(G_{j}^{\top}G_{j}\right)$ $=tr\left(\Sigma_{c}^{\frac{1}{2}}e_{j}\mu^{\top}\Sigma_{c}^{-\frac{1}{2}}\Sigma_{c}^{-\frac{1}{2}}\mu e_{j}^{\top}\Sigma_{c}^{\frac{1}{2}}\right)$
$=tr\left(e_{j}^{\top}\Sigma_{c}^{\frac{1}{2}}\Sigma_{c}^{\frac{1}{2}}e_{j}\mu^{\top}\Sigma_{c}^{-1}\mu\right)$
$=(\Sigma_{c})_{jj}\mu^{\top}\Sigma_{c}^{-1}\mu\leqslant V_{2}S_{2}\,,$
where the bound in last line follows Assumption 5.6 and Lemma A.1. Note that $P_{i}^{\top}G_{j}P_{i}$ (for $i$ with $A_{i}=0$) are i.i.d with mean $\mu$. Thus, an application of Hanson-Wright inequality (Theorem 1.1 of Rudelson and Vershynin, (2013)) leads to:
$\mathbb{E}\left[{\sf exp}\left\{t\left(P_{i}^{\top}G_{j}P_{i}-\mu_{j}\right)\right\}\mid A_{i}=0\right]\leqslant{\sf exp}\left\{C_{3}t^{2}\varsigma^{4}V_{2}S_{2}\right\},\quad\forall t\leqslant\frac{C_{4}}{\varsigma^{2}\sqrt{V_{2}S_{2}}}.$ (A.16)
This implies for all $t\leq\sqrt{n_{c}}C_{4}/(\varsigma^{2}\sqrt{V_{2}S_{2}})$, we obtain:
$\mathbb{E}\left[{\sf exp}\left\{\sqrt{n_{c}}t\left(\xi_{3}-\mu\right)_{j}\right\}\mid A_{1:n}\right]$ $=\mathbb{E}\left[{\sf exp}\left\{\frac{t}{\sqrt{n_{c}}}\sum_{\{i:A_{i}=0\}}(P_{i}^{\top}G_{j}P_{j}-\mu_{j})\right\}\mid A_{1:n}\right]$
=$=\prod_{\{i:A_{i}=0\}}\mathbb{E}\left[{\sf exp}\left\{\frac{t}{\sqrt{n_{c}}}(P_{i}^{\top}G_{j}P_{j}-\mu_{j})\right\}\mid A_{j}=0\right]$
$\leq\prod_{\{i:A_{i}=0\}}{\sf exp}\left\{C_{3}^{2}(t^{2}/n_{c})\varsigma^{4}V_{2}S_{2}\right\}$
$={\sf exp}\left\{C_{3}^{2}t^{2}\varsigma^{4}V_{2}S_{2}\right\}\,.$ (A.17)
Then, applying Chernoff’s inequality, we conclude
$\mathbb{P}\left(\sqrt{n_{c}}\|\xi_{3}-\mu\|_{\infty}\geq 2C_{3}^{2}t\varsigma^{4}V_{2}S_{2}\mid A_{1:n}\right)$ $\leq\sum_{j=1}^{p}\mathbb{P}\left(\sqrt{n_{c}}\left(\xi_{3}-\mu\right)_{j}\geqslant 2C_{3}^{2}t\varsigma^{4}V_{2}S_{2}\mid A_{1:n}\right)$
$=\sum_{j=1}^{p}\mathbb{P}\left(\sqrt{n_{c}}t\left(\xi_{3}-\mu\right)_{j}\geqslant 2C_{3}^{2}t^{2}\varsigma^{4}V_{2}S_{2}\mid A_{1:n}\right)$
$=\sum_{j=1}^{p}\mathbb{P}\left({\sf exp}{(\sqrt{n_{c}}t\left(\xi_{3}-\mu\right)_{j})}\geq{\sf exp}{(2C_{3}^{2}t^{2}\varsigma^{4}V_{2}S_{2})}\mid A_{1:n}\right)$
$\leq\sum_{j=1}^{p}{\sf exp}{(-2C_{3}^{2}t^{2}\varsigma^{4}V_{2}S_{2})}\mathbb{E}\left[{\sf exp}{(\sqrt{n_{c}}t\left(\xi_{3}-\mu\right)_{j})}\mid A_{1:n}\right]$
$={\sf exp}{(\log{(p)}-C_{3}^{2}t^{2}\varsigma^{4}V_{2}S_{2})}\,.$ (A.18)
The above inequality holds for any choice of $t\leq\sqrt{n_{c}}C_{3}/\varsigma^{2}\sqrt{V_{2}S_{2}}$. By selecting $t=\sqrt{2\log(p)}/\varsigma^{2}\sqrt{C_{3}^{2}V_{2}S_{2}}$, we have from Equation (A.2)
$\mathbb{P}\left(\sqrt{n_{t}}\|\xi_{3}-\mu\|_{\infty}\geq 2C_{3}^{2}t\varsigma^{4}V_{2}S_{2}\mid A_{1:n}\right)$ $=\mathbb{P}\left(\|\xi_{3}-\mu\|_{\infty}\geq 2C_{3}\varsigma^{2}\sqrt{2V_{2}S_{2}}\sqrt{\frac{\log{(p)}}{n_{c}}}\mid A_{1:n}\right)$
$\leq p^{-1}\,.$ (A.19)
Defining $K_{7}=C_{3}\varsigma^{2}\sqrt{2V_{2}s_{2}}$ and taking expectation with respect to to $\sigma$-field $\mathcal{F}_{0,n}$ on the both side of Equation (A.2), we conclude that $\|\xi_{3}-\mu\|_{\infty}\leq K_{7}\sqrt{\log(p)/n_{c}}$ with high probability.
We will next provide the upper bound on $\|\xi_{4}\|_{\infty}$. By the definition of $\xi_{4}$ (Equation (A.2)), we obtain:
$\|\xi_{4}\|_{\infty}$ $\leq\left\|\mu_{c}\right\|_{\infty}\left|\frac{1}{n_{c}}\sum_{\{i:A_{i}=0\}}(\mathbf{X}_{c,i}-\mu_{c})^{\top}\Sigma_{c}^{-1}\mu\right|$
$=\left\|\mu_{c}\right\|_{\infty}\left|\frac{1}{n_{c}}\sum_{\{i:A_{i}=0\}}P_{i}^{\top}\Sigma_{c}^{-1/2}\mu\right|$
$:=\left\|\mu_{c}\right\|_{\infty}\left|\frac{1}{n_{c}}\sum_{\{i:A_{i}=0\}}Z_{c,i}\right|\,,$
where $Z_{c,i}=P_{i}^{\top}\Sigma_{c}^{-1/2}\mu$. By Chebychev’s inequality, we obtain:
$\mathbb{P}\left(\left|\frac{1}{n_{c}}\sum_{\{i:A_{i}=0\}}Z_{c,i}\right|\geq t\mid A_{1:n}\right)\leq\frac{\operatorname{Var}\left(Z_{c}\mid A_{i}=0\right)}{n_{c}t^{2}}\,.$
Observe that $\operatorname{Var}(Z_{c}\mid A_{i}=0)=\mu^{\top}\Sigma_{c}^{-1}\mu$ which is upper bounded by $V_{2}$ by Lemma A.1. Therefore, choosing $t=\sqrt{\log{(p)}/n_{c}}$, we have:
$\mathbb{P}\left(\left|\frac{1}{n_{t}}\sum_{\{i:A_{i}=0\}}Z_{c,i}\right|\geq\sqrt{\frac{\log{(p)}}{n_{c}}}\mid A_{1:n}\right)\leq\frac{\operatorname{Var}\left(Z_{c}\mid A_{i}=0\right)}{\log{(p)}}\,.$
Moreover, $\|\mu_{c}\|_{\infty}\leq C_{c}$ by Assumption 5.6. Hence, we conclude that $\|\xi_{4}\|_{\infty}\leq K_{8}\sqrt{\log{(p)}/n_{c}}$ with probability going to 1. Finally, combining this bound along with the upper bounds in Equation (A.13) and (A.15) we conclude that $\|a-\mathbf{W}_{t}^{\top}\tau_{1}^{*}\|_{\infty}\leq K_{5}\sqrt{\log{(p)}/n_{c}}$ with probability going to 1, for some constant $K_{5}>0$.
Ultimately, we can derive the desired result $\|\tau_{2}^{*}\|_{\infty}\leqslant n_{c}^{-2/3}$ with probability tending to 1 under Assumption 5.6. This follows from applying the union bound, leveraging the sub-Gaussian properties of $P_{i}$, and mirroring the proof strategy used to establish $\|\tau_{1}^{*}\|_{\infty}\leq n_{t}^{-2/3}$. Thus, the proof is concluded. $\square$
Lemma A.1.
Under the Assumption 5.6, where $\tilde{a}^{\top}\Sigma_{t}^{-1}\tilde{a}=\mathcal{O}(1)$ and $\mu^{\top}\Sigma_{c}^{-1}\mu=\mathcal{O}(1)$, we have $a^{\top}\Sigma_{t}^{-1}a=\mathcal{O}_{p}(1)$ and $\bar{X}^{\top}\Sigma_{c}^{-1}\bar{X}=\mathcal{O}_{p}(1)$. Also, we have $\check{a}^{\top}\Sigma_{t}^{-1}\check{a}=\mathcal{O}_{p}(1)$ where $\check{a}=[\mu\quad\hat{\mathbf{B}}_{0}\mu]$, i.e., for some constant $V_{1}^{\prime}$, $\check{a}^{\top}\Sigma_{t}^{-1}\check{a}\leq V_{1}^{\prime}$ with high probability.
A.3 Proof of Lemma A.1
To prove the result of $a^{\top}\Sigma_{t}^{-1}a=\mathcal{O}_{p}(1)$ and $\bar{X}^{\top}\Sigma_{c}^{-1}\bar{X}=\mathcal{O}_{p}(1)$, where $a=[\bar{X}\quad\hat{\mathbf{B}}_{0}\bar{X}]$ , we will first focus on the behavior of $\bar{X}$ and $\hat{\mathbf{B}}_{0}$.
For any $\tilde{\lambda}\geq 4\sigma_{2}\sqrt{2S_{2}\log(pe^{t^{2}/2})/n_{c}}$, we establish
$\mathbb{P}\left(\max_{1\leqslant j\leqslant q}\left\|\left(\mathbf{B}_{0}-\hat{\mathbf{B}}_{0}\right)_{j}\right\|_{1}\geqslant t\right)\leqslant\sum_{j}\mathbb{P}\left(\left\|\left(\mathbf{B}_{0}-\hat{\mathbf{B}}_{0}\right)_{j}\right\|_{1}\geqslant\frac{4\tilde{\lambda}s}{\varphi_{c}^{2}}\right)\leqslant qp^{-\tilde{c}}\,,$
by selecting $e^{t^{2}/2}=p^{\tilde{c}}$, where $\varphi_{c}$ is a compatibility constant in Assumption 5.9. The last line follows the union bound, Theorem 6.1 and the generalization of Lemma 6.2 in Bühlmann and Van De Geer, (2011). (Here, we generalize Lemma 6.2 in Bühlmann and Van De Geer, (2011) to the scenario where $(\Sigma_{c})_{jj}\leqslant S_{2}$ rather than the original condition $(\Sigma_{c})_{jj}=1$. This relaxation remains valid by appropriately setting $S_{2}=1$.). Hence, by leveraging the sparsity Assumption 5.4 and Assumption 5.6, where $\frac{q}{p^{\tilde{c}}}\rightarrow 0$ as $q\rightarrow\infty$, or equivalently, there exists $\tilde{c}$ such that $q\leqslant p^{\tilde{c}}$, we establish the following bound:
$\left\|\left(\mathbf{B}_{0}-\mathbf{\hat{B}}_{0}\right)\right\|_{1,\infty}=\mathcal{O}_{p}\left(s\sqrt{\log(p)/n_{c}}\right)\,.$ (A.20)
Then, by applying Hoeffding’s inequality, we can deduce that, with high probability:
$\|\bar{X}-\mathbb{E}[X]\|_{\infty}\leqslant C_{X}\sqrt{\log(p)/n}.$ (A.21)
Consequently, we conclude that $a^{\top}\Sigma_{t}^{-1}a=\mathcal{O}_{p}(1)$ and $\bar{X}^{\top}\Sigma_{c}^{-1}\bar{X}=\mathcal{O}_{p}(1)$.
Similarity, as $\left\|\left(\mathbf{B}_{0}-\mathbf{\hat{B}}_{0}\right)\right\|_{1,\infty}=\mathcal{O}_{p}\left(s\sqrt{\log(p)/n_{c}}\right)$ we can conclude that for some constant $V_{1}^{\prime}$, $\check{a}^{\top}\Sigma_{t}^{-1}\check{a}\leq V_{1}^{\prime}$ with high probability. $\square$
A.4 Proof of Theorem 5.11
For notational convenience, in the following proof we drop the superscript of $U_{i}^{(0)}$ and $\epsilon_{i}^{(1)}$. Recall from the Proposition 4.1:
$\frac{\sqrt{n}(\hat{\theta}_{0}-\theta_{0})}{\sigma_{n}}$ $=\frac{\sqrt{n}}{\sigma_{n}}\left\{\left\|a-\mathbf{W}_{t}^{\top}\tau_{1}\right\|_{\infty}\left\|\hat{\phi}_{1}-\phi_{1}\right\|_{1}+\left\|\bar{X}-\mathbf{X}_{c}^{\top}\tau_{2}\right\|_{\infty}\|\hat{b}-b\|_{1}+\|\mathbf{B}_{0}\bar{X}-\hat{\mathbf{B}}_{0}\bar{X}\|_{\infty}\|\hat{\gamma}_{1}-\gamma_{1}\|_{1}\right\}$
$\qquad\qquad+\frac{\sqrt{n}}{\sigma_{n}}\left\{\sum_{\{i:A_{i}=1\}}\tau_{1,i}\epsilon_{i}+\sum_{\{i:A_{i}=0\}}\tau_{2,i}U_{i}^{\top}\hat{\gamma}_{1}\right\}$ (A.22)
$:=\frac{\sqrt{n}}{\sigma_{n}}\Delta_{n}+\frac{\sqrt{n}}{\sigma_{n}}V_{n}\,,$
where
$\Delta_{n}$ $=\left\|a-\mathbf{W}_{t}^{\top}\tau_{1}\right\|_{\infty}\left\|\hat{\phi}_{1}-\phi_{1}\right\|_{1}+\left\|\bar{X}-\mathbf{X}_{c}^{\top}\tau_{2}\right\|_{\infty}\|\hat{b}-b\|_{1}+\|\mathbf{B}_{0}\bar{X}-\hat{\mathbf{B}}_{0}\bar{X}\|_{\infty}\|\hat{\gamma}_{1}-\gamma_{1}\|_{1}\,,$
$V_{n}$ =$=\sum_{\{i:A_{i}=1\}}\tau_{1,i}\epsilon_{i}+\sum_{\{i:A_{i}=0\}}\tau_{2,i}U_{i}^{\top}\hat{\gamma}_{1}\,.$
If we can demonstrate that the error term $\sqrt{n}|\Delta_{n}|/\sigma_{n}=o_{p}(1)$ and that $\sqrt{n}V_{n}/\sigma_{n}$ converges to $\mathcal{N}(0,1)$, then we are done.
Asymptotic Negligibility of $\sqrt{n}\Delta_{n}/\sigma_{n}$: We establish this in two parts; we first show that $\sqrt{n}\Delta_{n}=o_{p}(1)$, and then we argue that $\sigma_{n}$ remains bounded away from 0. For the first step, observe that the standard estimation error of the Lasso estimator indicates:
$\|\hat{\phi}_{1}-\phi_{1}\|_{1}=\mathcal{O}_{p}\left(k\sqrt{\frac{\log(p+q)}{n_{t}}}\right),\quad\|\hat{b}-b\|_{1}=\mathcal{O}_{p}\left(sk\sqrt{\frac{\log(p)}{n_{c}}}\right),\quad\|\hat{\gamma}_{1}-\gamma_{1}\|_{1}=\mathcal{O}_{p}\left(k\sqrt{\frac{\log(q)}{n_{t}}}\right).$ (A.23)
The above risk bounds rely on the sub-gaussianity of the covariates and error terms (Assumption 5.9) and restricted eigenvalue (RE) condition of the sample covariance matrix. To obtain the rate of convergence of $\|\hat{\phi}_{1}-\phi_{1}\|_{1}$ and $\|\hat{\gamma}_{1}-\gamma_{1}\|_{1}$ we need $(\mathbf{W}_{t}^{\top}\mathbf{W}_{t})/n_{t}$ to satisfy the RE condition, and for obtaining the rate of convergence of $\|\hat{b}-b\|_{1}$, we need $(\mathbf{X}_{c}^{\top}\mathbf{X}_{c})/n_{c}$ to satisfy the RE condition. As their expected value is $\Sigma_{t}+\mu_{t}\mu_{t}^{\top}\succeq\Sigma_{t}$ and $\Sigma_{c}+\mu_{c}\mu_{c}^{\top}\succeq\Sigma_{c}$ and they satisfy the RE condition (Assumption 5.9), the sample second moment matrices $(\mathbf{W}_{t}^{\top}\mathbf{W}_{t})/n_{t}$, $(\mathbf{X}_{c}^{\top}\mathbf{X}_{c})/n_{c}$ also satisfy RE condition with high probability by Theorem 6 of Rudelson and Zhou, (2012). Next, we can use Corollary 2 of Negahban et al., (2012) for a column-standardized design matrix. Theorem 2.1 in Rudelson and Vershynin, (2013) can be applied to verify that our design matrices $\mathbf{W}_{t}$ and $\mathbf{X}_{c}$ are column-standardized with high probability,
$n_{t}^{-1/2}\|\mathbf{W}_{t,j}\|_{2}$ $\leq n_{t}^{-1/2}\|\mathbf{W}_{t,j}-\mu_{t}\|_{2}+\underbrace{\|\mu_{t}\|_{\infty}}_{\mathcal{O}(1)}\leq\left(\frac{c_{1}}{4}\iota^{2}\right)S_{1}^{1/2},\forall j=\{1,\ldots,p+q\},$ (A.24)
$n_{c}^{-1/2}\|\mathbf{X}_{c,k}\|_{2}$ $\leq n_{c}^{-1/2}\|\mathbf{X}_{c,k}-\mu_{c}\|_{2}+\underbrace{\|\mu_{c}\|_{\infty}}_{\mathcal{O}(1)}\leq\left(\frac{c_{2}}{4}\varsigma^{2}\right)S_{2}^{1/2},\forall k=\{1,\ldots,p\}.$
where the $\mathcal{O}(1)$ follows the Assumption 5.6. Hence, pairing these facts with the sparsity and sub-Gaussian noise assumptions in Assumptions 5.4 and 5.6, along with the condition stated in (A.24), Corollary 2 of Negahban et al., (2012) implies that $\hat{\phi}_{1}$, $\hat{b}$, and $\hat{\gamma}_{1}$, obtained by applying Lasso with parameters $(\lambda_{1},\lambda_{2})$ as specified in Theorem 5.11, satisfy the $\ell_{1}$-risk bound in Equation (A.23) with high probability.
The bounds for $\|a-\mathbf{W}_{t}^{\top}\tau_{1}\|_{\infty}$ and $\|\bar{X}-\mathbf{X}_{c}^{\top}\tau_{2}\|_{\infty}$ have already been established in the proof of Lemma 5.10. Next, we proceed to analyze the bound of the third product term, $\|\mathbf{B}_{0}\bar{X}-\hat{\mathbf{B}}_{0}\bar{X}\|_{\infty}\|\hat{\gamma}_{1}-\gamma_{1}\|_{1}$, in $\Delta_{n}$ as given in (A.4). Specifically, we write:
$\|(\mathbf{B}_{0}-\hat{\mathbf{B}}_{0})\bar{X}\|_{\infty}\|\hat{\gamma}_{1}-\gamma_{1}\|_{1}\leqslant$ $\max_{1\leqslant j\leq q}\left|\left\langle\left(\mathbf{B}_{0}-\hat{\mathbf{B}}_{0}\right)_{j},\bar{X}\right\rangle\right|\|\hat{\gamma}_{1}-\gamma_{1}\|_{1}$
$\leqslant$ $\max_{1\leqslant j\leq q}\|\bar{X}\|_{\infty}\left\|\left(\mathbf{B}_{0}-\hat{\mathbf{B}}_{0}\right)_{j}\right\|_{1}\|\hat{\gamma}_{1}-\gamma_{1}\|_{1}$
$\leqslant$ $\|\bar{X}\|_{\infty}\left[\max_{1\leqslant j\leqslant q}\left\|\left(\mathbf{B}_{0}-\hat{\mathbf{B}}_{0}\right)_{j}\right\|_{1}\right]\|\hat{\gamma}_{1}-\gamma_{1}\|_{1}$
$=$ $\|\bar{X}-\mu+\mu\|_{\infty}\left[\max_{1\leqslant j\leqslant q}\left\|\left(\mathbf{B}_{0}-\hat{\mathbf{B}}_{0}\right)_{j}\right\|_{1}\right]\|\hat{\gamma}_{1}-\gamma_{1}\|_{1}$
$\leqslant$ $(\underbrace{\|\bar{X}-\mu\|_{\infty}}_{o_{p}(1)}+\underbrace{\|\mu\|_{\infty}}_{\leqslant C_{X}})\left[\max_{1\leqslant j\leqslant q}\left\|\left(\mathbf{B}_{0}-\hat{\mathbf{B}}_{0}\right)_{j}\right\|_{1}\right]\|\hat{\gamma}_{1}-\gamma_{1}\|_{1}$
$\leqslant$ $2C_{X}\left[\max_{1\leqslant j\leqslant q}\left\|\left(\mathbf{B}_{0}-\hat{\mathbf{B}}_{0}\right)_{j}\right\|_{1}\right]\|\hat{\gamma}_{1}-\gamma_{1}\|_{1}$
$=$ $2C_{X}\left\|\mathbf{B}_{0}-\hat{\mathbf{B}}_{0}\right\|_{1,\infty}\left\|\hat{\gamma}_{1}-\gamma_{1}\right\|_{1}\,,$
where $\|\mathbf{B}_{0}-\hat{\mathbf{B}}_{0}\|_{1,\infty}=\mathcal{O}_{p}(s\sqrt{\log(p)/n_{c}})$ as shown in Equation (A.20) from Lemma A.1. Therefore, we can analyze the bound of $\Delta_{n}$ separately as follows: from Equations (A.23) and (A.2), we derive the upper bounds $\|a-\mathbf{W}_{t}^{\top}\tau_{1}\|_{\infty}=\mathcal{O}_{p}(\sqrt{\log(p+q)/n_{t}})$ and $\|\hat{\phi}_{1}-\phi_{1}\|_{1}=\mathcal{O}_{p}(k\sqrt{\log(p+q)/n_{t}})$. By combining these two bounds, we obtain:
$\left\|a-\mathbf{W}_{t}^{\top}\tau_{1}\right\|_{\infty}\left\|\hat{\phi}_{1}-\phi_{1}\right\|_{1}=\mathcal{O}_{p}\left(k\frac{\log(p+q)}{n_{t}}\right)\,.$ (A.25)
Similarly, by combining the upper bounds from Equations (A.23) and (A.12), where $\|\bar{X}-\mathbf{X}_{c}^{\top}\tau_{2}\|_{\infty}=\mathcal{O}_{p}\left(\sqrt{\log(p)/n_{c}}\right)$ and $\|\hat{b}-b\|_{1}=\mathcal{O}_{p}\left(sk\sqrt{\log(p)/n_{c}}\right)$, we obtain:
$\left\|\bar{X}-\mathbf{X}_{c}^{\top}\tau_{2}\right\|_{\infty}\|\hat{b}-b\|_{1}=\mathcal{O}_{p}\left(sk\frac{\log(p)}{n_{c}}\right)\,.$ (A.26)
Combining the upper bound in Equation (A.20) and (A.23), $\|(\mathbf{B}_{0}-\hat{\mathbf{B}}_{0})\bar{X}\|_{\infty}=\mathcal{O}_{p}\left(s\sqrt{\log(p)/n_{c}}\right)$ and $\|\gamma_{1}-\hat{\gamma}_{1}\|_{1}=\mathcal{O}_{p}\left(k\sqrt{\log(q)/n_{t}}\right)$, we have:
$\|(\mathbf{B}_{0}-\hat{\mathbf{B}}_{0})\bar{X}\|_{\infty}\|\hat{\gamma}_{1}-\gamma_{1}\|_{1}=\mathcal{O}_{p}\left(sk\frac{\sqrt{\log(q)\log(p)}}{n_{t}\wedge n_{c}}\right)\,.$ (A.27)
Combining Equation (A.25), (A.26) together, along with the sparsity Assumption 5.4, we conclude that:
$\sqrt{n}\left\|a-\mathbf{W}_{t}^{\top}\tau_{1}\right\|_{\infty}\left\|\hat{\phi}_{1}-\phi_{1}\right\|_{1}+\sqrt{n}\left\|\bar{X}-\mathbf{X}_{c}^{\top}\tau_{2}\right\|_{\infty}\|\hat{b}-b\|_{1}+\sqrt{n}\|\mathbf{B}_{0}\bar{X}-\hat{\mathbf{B}}_{0}\bar{X}\|_{\infty}\|\hat{\gamma}_{1}-\gamma_{1}\|_{1}$
$=\mathcal{O}_{p}\left(k\frac{\log(p+q)}{\sqrt{n_{t}}}\right)+\mathcal{O}_{p}\left(sk\frac{\log(p)}{\sqrt{n_{c}}}\right)+\mathcal{O}_{p}\left(sk\frac{\sqrt{\log(q)\log(p)}}{\sqrt{n_{c}\wedge n_{t}}}\right)$
$=o_{p}(1)\,,$ (A.28)
where we use $n\sim n_{t}\sim n_{c}$. This concludes that $|\Delta_{n}|=o_{p}(n^{-1/2})$.
Our next goal is to show that $\sigma_{n}$ (or equivalently $\sigma_{n}^{2}$) remains bounded away from 0. Recall the definition of $\sigma_{n}^{2}$:
$\sigma_{n}^{2}=n\left(\sigma_{1}^{2}\|\tau_{1}\|_{2}^{2}+\sigma_{2}^{2}\|\tau_{2}\|_{2}^{2}\|\gamma_{1}\|_{2}^{2}\right)\,.$
Observe that the minimum estimand size on $a$ and $\bar{X}$ satisfies $\|a\|_{\infty}\geq 2\kappa$ and $\|\bar{X}\|_{\infty}\geq 2\kappa$ by application of triangle inequality and the results in (A.20) and (A.21). Therefore, any feasible solution to (5.4) and (5.5) should satisfy:
$|(\mathbf{W}_{t}^{\top}\tau_{1})_{j}|$ $\geq|a_{j}|-|a_{j}-(\mathbf{W}_{t}^{\top}\tau_{1})_{j}|\geq 2\kappa-K_{1}\sqrt{\frac{\log{(p+q)}}{n_{t}}}\geq\kappa\,,$
$|(\mathbf{X}_{c}^{\top}\tau_{2})_{j}|$ $\geq|\bar{X}_{j}|-|\bar{X}_{j}-(\mathbf{X}_{c}^{\top}\tau_{2})_{j}|\geq 2\kappa-K_{2}\sqrt{\frac{\log{p}}{n_{c}}}\geq\kappa\,,$
as $\log{(p+q)}/(n_{t}\wedge n_{c})\to 0$. Applying Cauchy-Schwarz inequality, we obtain:
$\kappa^{2}$ $\leq(\mathbf{W}_{t}^{\top}\tau_{1})_{j}^{2}\leq\|\tau_{1}\|_{2}^{2}\sum_{\{i:A_{i}=1\}}\mathbf{W}_{t,ij}^{2}\implies n\|\tau_{1}\|_{2}^{2}\geq\frac{\kappa^{2}}{\frac{n_{t}}{n}\frac{1}{n_{t}}\sum_{i:A_{i}=1}\mathbf{W}^{2}_{ij}}\,,$
$\kappa^{2}$ $\leq(\mathbf{X}_{c}^{\top}\tau_{2})_{j}^{2}\leq\|\tau_{2}\|_{2}^{2}\sum_{\{i:A_{i}=0\}}\mathbf{X}_{c,ij}^{2}\implies n\|\tau_{2}\|_{2}^{2}\geq\frac{\kappa^{2}}{\frac{n_{c}}{n}\frac{1}{n_{c}}\sum_{i:A_{i}=0}\mathbf{X}^{2}_{ij}}\,.$
These inequalities, along with the upper bound established in Equation (A.24), conclude that with probability approaching 1:
$n\|\tau_{1}\|_{2}^{2}\geq\frac{\kappa^{2}}{C_{1}},\ \ \ \ n\|\tau_{2}\|_{2}^{2}\geq\frac{\kappa^{2}}{C_{2}}\,.$
for some constant $C_{1},C_{2}>0$. Hence with probability going to 1:
$\sigma_{n}^{2}=n\left(\sigma_{1}^{2}\|\tau_{1}\|_{2}^{2}+\sigma_{2}^{2}\|\tau_{2}\|_{2}^{2}\|\gamma_{1}\|_{2}^{2}\right)\geq n\left(\sigma_{1}^{2}\|\tau_{1}\|_{2}^{2}+\sigma_{2}^{2}\|\tau_{2}\|_{2}^{2}\right)\geq\kappa^{2}\left(\frac{\sigma_{1}^{2}}{C_{1}}+\frac{\sigma_{2}^{2}}{C_{2}}\right)\,.$ (A.29)
This concludes the proof of the asymptotic ineligibility of the bias term.
Asymptotic normality of $\sqrt{n}V_{n}/\sigma_{n}$: Recall the definition of $V_{n}$:
$\frac{\sqrt{n}}{\sigma_{n}}V_{n}$ $:=\frac{\sqrt{n}}{\sigma_{n}}\left(\sum_{\left\{i:A_{i}=1\right\}}\tau_{1,i}\epsilon_{i}+\sum_{\{i:A_{i}=0\}}\tau_{2,i}U_{i}^{\top}\hat{\gamma}_{1}\right)$
$=\frac{\sqrt{n}}{\sigma_{n}}\left(\sum_{\left\{i:A_{i}=1\right\}}\tau_{1,i}\epsilon_{i}+\sum_{\{i:A_{i}=0\}}\tau_{2,i}U_{i}^{\top}\gamma_{1}\right)+\frac{\sqrt{n}}{\sigma_{n}}\left(\sum_{\{i:A_{i}=0\}}\tau_{2,i}U_{i}^{\top}(\hat{\gamma}_{1}-\gamma_{1})\right)$
$:=T_{1}+T_{2}\,.$
In the rest of the analysis, we show $T_{1}$ converges to $\mathcal{N}(0,1)$ and $T_{2}=o_{p}(1)$. Note that the expression of $T_{1}$ can be simplified to:
$T_{1}=\frac{1}{\sqrt{n(\sigma_{1}^{2}\|\tau_{1}\|_{2}^{2}+\sigma_{2}^{2}\|\tau_{2}\|_{2}^{2}\|\gamma_{1}\|_{2}^{2})}}\left(\sum_{\left\{i:A_{i}=1\right\}}\tau_{1,i}\epsilon_{i}+\sum_{\left\{i:A_{i}=0\right\}}\tau_{2,i}U_{i}^{\top}\gamma_{1}\right)\,,$
which follows by cancelling $\sqrt{n}$ from the definition of $\sigma_{n}$. Since the noise is centered i.i.d sub-Gaussian random variables (Assumption 5.5), we have conditional on $\mathcal{G}_{n}=\sigma(\mathcal{D}_{1},\mathbf{X},\mathbf{M}_{c},A_{1:n})$, $T_{1}$ has mean 0 and variance 1. Furthermore, it is easy to check that Lyapunov’s condition is also satisfied with $\delta=1$ as $\sqrt{n}\|\tau_{1}\|_{\infty}+\sqrt{n}\|\tau_{2}\|_{\infty}=o_{p}(1)$ and $\|\gamma_{1}\|_{2}$ is bounded. Therefore, an application of the conditional central limit theorem (Bulinski, (2017)) yields:
$\mathbb{P}(T_{1}\leq t\mid\mathcal{G}_{n})\longrightarrow e^{-\frac{t^{2}}{2}},\ \ \ \text{a.s. in }(X_{1:n},M_{1:n},A_{1:n})\,.$
Finally, taking the expectation with respect to $\mathcal{G}_{n}$ and applying dominated convergence theorem (as $\mathbb{P}(T_{1}\leq t\mid\mathcal{G}_{n})\leq 1$), we conclude that $T_{1}$ converges to $\mathcal{N}(0,1)$.
To conclude the proof, all we need to show is $T_{2}=o_{p}(1)$. As we have already proved, $\sigma_{n}$ stays bounded away from 0 with probability approaching 1 (Equation (A.29)), all we need to show is that $\sqrt{n}\sum_{\{i:A_{i}=0\}}\tau_{2,i}U_{i}^{\top}(\hat{\gamma}_{1}-\gamma_{1})=o_{p}(1)$. In the rest of the analysis, we use the fact that by definition of the optimization problem Equation (5.4), $\|\tau_{2}\|_{2}^{2}\leq\|\tau_{2}^{*}\|_{2}^{2}$ whenever $\tau_{2}^{*}$ is feasible. In the proof of Lemma 5.10, we have proved that $\tau_{2}^{*}$ is feasible with probability going to 1. Call this event, i.e., $\tau_{2}^{*}$ is a feasible solution of the optimization problem (5.4), $\Omega_{n}$. Therefore, i) $\mathbb{P}(\Omega_{n}^{c})\to 0$ as $n\uparrow\infty$, and ii) On $\Omega_{n}$, we have $\|\tau_{2}\|_{2}^{2}\leq\|\tau_{2}^{*}\|_{2}^{2}$. Now we have:
$\mathbb{P}\left(\sqrt{n}\sum_{\{i:A_{i}=0\}}\tau_{2,i}U_{i}^{\top}(\hat{\gamma}_{1}-\gamma_{1})\geq t\right)$
$=\mathbb{P}\left(\sqrt{n}\sum_{\{i:A_{i}=0\}}\tau_{2,i}U_{i}^{\top}(\hat{\gamma}_{1}-\gamma_{1})\geq t,\Omega_{n}\right)+\mathbb{P}\left(\sqrt{n}\sum_{\{i:A_{i}=0\}}\tau_{2,i}U_{i}^{\top}(\hat{\gamma}_{1}-\gamma_{1})\geq t,\Omega_{n}^{c}\right)$
$\leq\mathbb{P}\left(\sqrt{n}\sum_{\{i:A_{i}=0\}}\tau_{2,i}U_{i}^{\top}(\hat{\gamma}_{1}-\gamma_{1})\geq t,\Omega_{n}\right)+\mathbb{P}(\Omega_{n}^{c})\,.$
Therefore, we need to show that the first term of the above equation goes to 0 as $n\uparrow\infty$. For simplicity of presentation, define $\rho_{i}:=U_{i}^{\top}(\hat{\gamma}_{1}-\gamma_{1})$. Therefore, conditional of $\tilde{\mathcal{G}}_{n}=\sigma(\mathbf{X},\mathbf{M}_{c},A_{1:n},\{{\epsilon}_{i}:A_{i}=1\})$, the terms $\tau_{2,i}\rho_{i}$’s are independent. Therefore, an application of Chebychev’s inequality yields:
$\mathbb{P}\left(\left|\sqrt{n}\sum_{\{i:A_{i}=0\}}\tau_{2,i}\rho_{i}\right|\geq t\mid\tilde{\mathcal{G}}_{n}\right)\mathds{1}_{\Omega_{n}}$ $\leq\frac{n_{c}\operatorname{Var}(\sum_{\{i:A_{i}-0\}}\tau_{2,i}\rho_{i}\mid\tilde{\mathcal{G}}_{n})}{t^{2}}\mathds{1}_{\Omega_{n}}$
$\leq\frac{n_{c}\sum_{\{i:A_{i}-0\}}\tau_{2,i}^{2}\operatorname{Var}(\rho_{i}\mid\tilde{\mathcal{G}}_{n})}{t^{2}}\mathds{1}_{\Omega_{n}}$
$\leq\frac{n_{c}\sigma_{2}^{2}\|\hat{\gamma}_{1}-\gamma_{1}\|_{2}^{2}}{t^{2}}\left(\sum_{\{i:A_{i}=0\}}\tau_{2,i}^{2}\right)\mathds{1}_{\Omega_{n}}$
$\leq\frac{n_{c}\sigma_{2}^{2}\|\hat{\gamma}_{1}-\gamma_{1}\|_{2}^{2}}{t^{2}}\left(\sum_{\{i:A_{i}=0\}}(\tau^{*}_{2,i})^{2}\right)\mathds{1}_{\Omega_{n}}\,.$ (A.30)
We now the definition of $\tau^{*}_{2,i}$ as defined in the proof of Lemma A.1.
$\sum_{\{i:A_{i}=0\}}\tau_{2,i}^{*2}=\sum_{\{i:A_{i}=0\}}\left(\frac{1}{n_{c}}\mu^{\top}\Sigma_{c}^{-1}(X_{c,i}-\mu_{c})\right)^{2}$ $=\frac{1}{n_{c}^{2}}\sum_{\{i:A_{i}=0\}}\left(\mu^{\top}\Sigma_{c}^{-1/2}P_{i}\right)^{2}$
$=\frac{1}{n_{c}^{2}}\sum_{\{i:A_{i}=0\}}\mu^{\top}\Sigma_{c}^{-1/2}P_{i}P_{i}^{\top}\Sigma_{c}^{-1/2}\mu$
$=\frac{1}{n_{c}}\mu^{\top}\Sigma_{c}^{-1/2}\left(\frac{1}{n_{c}}\sum_{\{i:A_{i}=0\}}P_{i}P_{i}^{\top}\right)\Sigma_{c}^{-1/2}\mu$
$=\frac{1}{n_{c}}d^{\top}\left(\frac{1}{n_{c}}\sum_{\{i:A_{i}=0\}}P_{i}P_{i}^{\top}\right)d$
$=\frac{1}{n_{c}}\left(\frac{1}{n_{c}}\sum_{\{i:A_{i}=0\}}Z_{i}^{2}\right)\,,$ (A.31)
where $d=\Sigma_{c}^{-1/2}\mu$ and $Z_{i}=P_{i}^{\top}d$ where $Z_{i}$’s are mean 0 and independent. Therefore, we have:
$\mathbb{P}\left(\left|\sqrt{n}\sum_{\{i:A_{i}=0\}}\tau_{2,i}\rho_{i}\right|\geq t\mid\tilde{\mathcal{G}}_{n}\right)\mathds{1}_{\Omega_{n}}\leq\frac{\sigma_{2}^{2}\|\hat{\gamma}_{1}-\gamma_{1}\|_{2}^{2}}{t^{2}}\left(\frac{1}{n_{c}}\sum_{\{i:A_{i}=0\}}Z_{i}^{2}\right)\,.$ (A.32)
Taking expectation on the both side, we have:
$\mathbb{P}\left(\left|\sqrt{n}\sum_{\{i:A_{i}=0\}}\tau_{2,i}\rho_{i}\right|\geq t,\Omega_{n}\right)$
$=\mathbb{E}\left[\mathbb{P}\left(\left|\sqrt{n}\sum_{\{i:A_{i}=0\}}\tau_{2,i}\rho_{i}\right|\geq t\mid\tilde{\mathcal{G}}_{n}\right)\mathds{1}_{\Omega_{n}}\right]$
$\leq\frac{\sigma_{2}^{2}}{t^{2}}\mathbb{E}\left[\|\hat{\gamma}_{1}-\gamma_{1}\|_{2}^{2}\left(\frac{1}{n_{c}}\sum_{\{i:A_{i}=0\}}Z_{i}^{2}\right)\right]$
$=\frac{\sigma_{2}^{2}}{t^{2}}\mathbb{E}[\|\hat{\gamma}_{1}-\gamma_{1}\|_{2}^{2}]\mathbb{E}\left[\frac{1}{n_{c}}\sum_{\{i:A_{i}=0\}}Z_{i}^{2}\right]\hskip 36.135pt[\text{As $\hat{\gamma}_{1}$ is obtained from treatment observations}]$
$\leq\frac{\sigma_{2}^{2}}{t^{2}}\mathbb{E}[\|\hat{\gamma}_{1}-\gamma_{1}\|_{2}^{2}](\mu^{\top}\Sigma_{c}^{-1}\mu)$
[…]
We conclude the proof by noting that $\mathbb{E}[\|\hat{\gamma}_{1}-\gamma_{1}\|_{2}^{2}]=o(1)$. $\square$
We present here for the ease of the presentation for Hanson-Wright type of inequality for quadratic forms in Sub-Gaussian non-necessarily independent random variables (Zajkowski, , 2020).
Lemma A.2.
[…]
$\mathbb{P}\left(|X^{\top}\mathbf{A}X-\mathbb{E}[X^{\top}\mathbf{A}X]|>t\right)\leq 2{\sf exp}\left(-\min\left\{\frac{t^{2}}{C_{3}^{2}\|\mathbf{A}\|_{F}\mathcal{K}^{4}},\frac{t}{C_{3}\|\mathbf{A}\|_{F}\mathcal{K}^{2}}\right\}\right)\,,$ (A.33)
where $C_{3}=2\sqrt{2}C_{1}C_{2}$.